### v2

# Reflected generalized BDSDEs driven by non-homogeneous Lévy processes and obstacle problems for stochastic integro-PDEs with nonlinear Neumann boundary conditions

Abstract
We consider reflected generalized backward doubly stochastic differential equations driven by a non-homogeneous Lévy process. Under stochastic conditions on the coefficients, we prove the existence and uniqueness of a solution. Furthermore, we apply these results to obtain a probabilistic representation for the viscosity solutions of an obstacle problem governed by stochastic integro-partial differential equations with a nonlinear Neumann boundary condition.
Keywords: Reflected generalized backward doubly SDEs; Non-homogeneous Lévy process; Obstacle; stochastic monotone condition; SIPDE; Viscosity solution; Nonlinear Neumann boundary conditions.
MSC 2020: 60H05; 60H15; 60H20; 60H30.
1 Introduction
Backward stochastic differential equations (BSDEs) were first introduced as adjoint equations in stochastic control and studied in depth by Bismut [3]. Later, Pardoux and Peng [35] established existence and uniqueness for nonlinear BSDEs. Since then, BSDEs have been widely applied in finance [10], insurance [5], risk management [13], stochastic optimal control and differential games [17, 18], and in the analysis of partial differential equations (PDEs) [33, 34, 39, 40, 41], where BSDEs are naturally linked to viscosity solutions of certain quasi-linear parabolic PDEs via probabilistic representations that extend the classical Feynman–Kac formula for linear parabolic PDEs.
Motivated by this latter connection, Pardoux and Zhang [38] introduced a new class of BSDEs that includes an additional finite-variation term, given by an integral with respect to a continuous increasing process interpreted as the boundary local time of a diffusion. These are called generalized BSDEs. This class has proved to be a powerful tool for providing probabilistic representations of solutions to systems of parabolic and elliptic semilinear PDEs with nonlinear Neumann boundary conditions. A solution on a time horizon $T\in(0,+\infty)$ with terminal value $\xi$ and drivers $f(\omega,t,y,z)$ and $h(\omega,t,y)$ is a pair of processes $(Y_{t},Z_{t})_{t\leq T}$ satisfying
$Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}-\int_{t}^{T}Z_{s}dW_{s},$ (1.1)
a.s. for all $t\in[0,T]$, where $W$ is a Brownian motion, $(\kappa_{t})_{t\geq 0}$ is a continuous increasing process with $\kappa_{0}=0$ (the local time), and $(Y,Z,\kappa)$ are adapted to the natural filtration of $W$. This framework was later extended by Pardoux [32], who considered a generalized BSDE driven by a Brownian motion together with independent Poisson jumps, and by Elmansouri and El Otmani [14, 15] to more general filtrations.
The generalized BSDE (1.1) was further extended to the reflected case by Ren and Xia [44]. Here, the state process $(Y_{t})_{t\leq T}$ is constrained to stay above a given barrier (obstacle) $S:=(S_{t})_{t\leq T}$, a continuous adapted process, following the formulation initiated by El Karoui et al. [9] for classical reflected BSDEs (the case $d\kappa=0$). The solution is now a triple $(Y_{t},Z_{t},K_{t})_{t\leq T}$ of adapted processes solving
$\left\{\begin{split}&~Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}+\left(K_{T}-K_{t}\right)-\int_{t}^{T}ZdW_{s},\quad t\in[0,T],\\ &~Y_{t}\geq S_{t},~t\in[0,T],\quad\text{ and }\int_{0}^{T}\left(Y_{s}-S_{s}\right)dK_{s}=0.\end{split}\right$. (1.2)
The main motivation for studying the reflected generalized BSDE (1.2) lies in its applications to optimal stopping and in providing a probabilistic representation of the viscosity solution to obstacle problems for PDEs with nonlinear Neumann boundary conditions. This line of research has attracted considerable attention; in particular, Ren and El Otmani [43] treated the reflected generalized BSDE driven by a Lévy process, and Elhachemy and El Otmani [12] considered the Brownian–Poisson setting, generalizing (1.2) by incorporating jumps together with stochastically monotone generators. In both works [12, 43], existence and uniqueness were applied to derive probabilistic formulas for the viscosity solution of obstacle problems for partial differential–integral equations with nonlinear Neumann boundary conditions (see, e.g., [2, 19]).
Beyond the (deterministic) PDE setting, an important direction is the probabilistic representation of solutions to stochastic PDEs (SPDEs), which have become a major focus of research and a key modeling tool in the applied sciences. The linear case was initiated by Pardoux [30, 31], who extended the Feynman–Kac formula and established basic existence and uniqueness results for SPDEs (see also [29, 46]). Quasilinear parabolic SPDEs were studied by Pardoux and Peng [36] through the theory of BSDEs. In 1994, Pardoux and Peng [36] introduced backward doubly stochastic differential equations (BDSDEs) as a generalization of (1.1), driven by both forward and backward Itô integrals. A solution is a pair $(Y_{t},Z_{t})_{t\leq T}$ such that
$Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}-\int_{t}^{T}Z_{s}dW_{s},\quad t\in[0,T].$ (1.3)
Here, $(W_{t})_{t\leq T}$ and $(B_{t})_{t\leq T}$ are independent Brownian motions; the integral with respect to $dW$ is the classical forward Itô integral, while the integral with respect to $\overleftarrow{dB}$ is a backward Itô integral. Within the framework of Nualart and Pardoux [25], these integrals are instances of the Itô Skorohod integral. In [36], existence and uniqueness for (1.3) are established under Lipschitz conditions on $f$ and $g$ together with appropriate square-integrability assumptions, and (1.3) is related to a probabilistic representation for a system of quasilinear SPDEs, thus extending [30, 31].
Motivated by Pardoux & Zhang [38] (for (1.2)) and Pardoux & Peng [36] (for (1.3)) and their applications to PDEs with nonlinear Neumann boundary conditions and to SPDEs, Boufoussi et al. [4] studied generalized BDSDEs. They proved existence and uniqueness under Lipschitz drivers and provided a probabilistic representation of stochastic viscosity solutions for semilinear SPDEs with a Neumann boundary condition.
A central ingredient in the BSDE theory is the martingale representation theorem. This result is classical in filtrations generated by a Brownian motion, a Poisson point process, or a Poisson random measure. For (homogeneous) Lévy processes, Nualart and Schoutens [26] established a martingale representation via the chaos expansion; building on this, [28] proved existence and uniqueness for BSDEs driven by Teugels martingales associated with Lévy processes possessing moments of all orders. El Otmani [11] studied generalized BSDEs driven by Teugels martingales together with an independent Brownian motion, while Hu and Ren [20] considered generalized BDSDEs driven by Lévy processes and provided a probabilistic interpretation for solutions to a class of stochastic integro-partial differential equations (SIPDEs) with a nonlinear Neumann boundary condition. Aman [1] treated the reflected case for such generalized BDSDEs and gave a probabilistic representation for reflected SIPDEs with a nonlinear Neumann boundary condition. More recently, El Jamali and El Otmani [8] extended the predictable representation theorem to non-homogeneous Lévy processes, using techniques akin to those of Nualart and Schoutens [26] in the homogeneous case. As an application, they studied BSDEs driven by a non-homogeneous Lévy process with a finite set of jump sizes and a driver with a stochastic Lipschitz coefficient $f$, of the form
$Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s},\quad t\in[0,T],$ (1.4)
where $(H^{(k)})_{k\geq 1}$ is a family of martingales, called the orthonormalized power-jump processes, associated with the components of the non-homogeneous Lévy process; see also their connection with IPDEs. For the generalized BSDE version of (1.4), El Jamali [7] proved existence and uniqueness under stochastic Lipschitz assumptions on the drivers. More recently, the doubly stochastic case of (1.4) was studied by Marzougue and Elmansouri [24], who considered the BDSDE
$Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}g(s,Y_{s})\overleftarrow{dB}_{s}-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s},\quad t\in[0,T],$
and established existence and uniqueness for the pair $(Y_{t},Z_{t})_{t\leq T}$ under stochastic Lipschitz conditions on $f$ and $g$. They also proved a comparison theorem and applied it to obtain existence and uniqueness in the general case where $f$ is continuous with stochastic linear growth, while $g$ remains stochastically Lipschitz.
Motivated by the above contributions, we study in this paper a reflected generalized backward doubly SDE driven by a non-homogeneous Lévy process (RGBDSDE-NL) of the form
$\left\{\begin{split}\text{(i).}&~Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}+\int_{t}^{T}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}+\left(K_{T}-K_{t}\right)\\ &\hskip 28.45274pt-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s},\quad t\in[0,T],\\ \text{(ii).}&~Y_{t}\geq S_{t},~t\in[0,T],\\ \text{(iii).}&\mbox{ Skorokhod condition: }\int_{0}^{T}\left(Y_{s}-S_{s}\right)dK_{s}=0.\end{split}\right$. (1.5)
Here,
  • the barrier $S:=(S_{t})_{t\leq T}$ is a continuous process with $S_{t}$ being $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$, where
    $\mathcal{F}_{t}\triangleq\mathcal{F}_{t}^{L}\otimes\mathcal{F}_{t,T}^{B}\vee\mathcal{N},\quad t\in[0,T],$
    $(B_{t})_{t\leq T}$ is a one-dimensional Brownian motion, $(L_{t})_{t\leq T}$ is a non-homogeneous Lévy process independent of $(B_{t})_{t\leq T}$, and $\mathcal{F}_{s,t}^{B}:=\sigma\{B_{r}-B_{s}:s\leq r\leq t\}$, while $\mathcal{F}_{t}^{L}:=\mathcal{F}_{0,t}^{L}$, and $\mathcal{N}$ is the collection of $\mathbb{P}$-null sets.
  • the drivers $f$ and $h$ satisfy stochastic monotonicity conditions with respect to the state variable $y$;
  • the drivers $f$ and $g$ are stochastically Lipschitz with respect to $z$ and $y$, respectively, and $g$ also satisfies a Lipschitz condition in $z$ as in [36];
  • $(\kappa_{t})_{t\leq T}$ is a continuous increasing process interpreted as the boundary local time of a Markov process driven by a non-homogeneous Lévy noise.
Under these general assumptions on the data, we establish existence and uniqueness for the RGBDSDE-NL (1.5). This extends and unifies the works [1, 4, 7, 20, 36] by encompassing previously studied deterministic Lipschitz and monotonicity conditions on the drivers. A second motivation of this work is the connection with SIPDEs. In particular, using the RGBDSDE-NL (1.5), we derive a probabilistic representation of the stochastic viscosity solution to an obstacle problem for stochastic integro-partial differential equations with nonlinear Neumann boundary conditions (SIPDE-NBC) of parabolic type, under suitable assumptions. We note that [1, 20] applied related results to SIPDE-NBCs with a (homogeneous) Lévy process without providing full details. In the non-homogeneous Lévy setting with stochastic monotonicity and Lipschitz conditions, the problem is new and technically challenging.
The paper is organized as follows. In Section 2, we specify the stochastic basis, introduce notations, and prove auxiliary results used throughout the paper. In Section 3, we state our assumptions on the data $(\xi,f,h,\kappa,g,S)$ of the RGBDSDE-NL (1.5), present a modified Itô formula for processes involving forward–backward integrals, and establish existence and uniqueness under our general stochastic conditions. Finally, in Section 4, we provide a probabilistic interpretation for solutions to a class of stochastic integro-partial differential equations with nonlinear Neumann boundary conditions.
2 Preliminaries
Let $T\in(0,+\infty)$ be a fixed horizon time. Throughout this paper, $(B_{t})_{t\leq T}$ will denote a standard one-dimensional Brownian motion and $(L_{t})_{t\leq T}$ a non-homogeneous Lévy process independent of $(B_{t})_{t\leq T}$ defined on the complete probability spaces $(\Omega_{1},\mathcal{F}_{1},\mathbb{P}_{1})$ and $(\Omega_{2},\mathcal{F}_{2},\mathbb{P}_{2})$, respectively.
We will denote by $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F}=\{\mathcal{F}_{t}\}_{0\leq t\leq T})$ the complete probability space, where $\Omega=\Omega_{1}\times\Omega_{2}$ $\mathcal{F}=\mathcal{F}_{1}\otimes\mathcal{F}_{2}$, $\mathbb{P}=\mathbb{P}_{1}\otimes\mathbb{P}_{2}$ and $\mathcal{F}_{t}\triangleq\mathcal{F}_{t}^{L}\otimes\mathcal{F}_{t,T}^{B}\vee\mathcal{N}$, where for any process $(\eta_{t})_{t\leq T}$; $\mathcal{F}_{s,t}^{\eta}=\sigma\{\eta_{r}-\eta_{s};\;s\leq r\leq t\}$, $\mathcal{F}_{t}^{\eta}:=\mathcal{F}_{0,t}^{\eta}$, $\mathbb{F}^{\eta}:=(\mathcal{F}_{t}^{\eta})_{t\leq T}$ and $\mathcal{N}$ is the collection of $\mathbb{P}$-null sets of $\mathcal{F}$. Note that the family of $\sigma$-algebras $\{\mathcal{F}_{t}\}_{t\leq T}$ is neither increasing nor decreasing; in particular, it is not a filtration. Any random variable $\zeta(\omega_{1})$, $\omega_{1}\in\Omega_{1}$ (resp. $\eta(\omega_{2})$, $\omega_{2}\in\Omega_{2}$) is considered as a random variable on $\Omega$ via the following identification: $\zeta(\omega)=\zeta(\omega_{1})$ (resp. $\eta(\omega)=\eta(\omega_{2})$), $\omega\triangleq(\omega_{1},\omega_{2})$.
By $\mathcal{B}(\mathbb{R})$, we mean the Borel $\sigma$-algebra of the real line $\mathbb{R}=(-\infty,+\infty)$. All stochastic processes are defined on the time interval $[0,T]$. Additionally, to simplify the notation, we omit the dependence on $\omega$ for any given process or random function. By convention, all brackets and stochastic integrals are assumed to be zero at time zero.
Definition 1.
An $\mathbb{F}^{L}$-adapted $\mathbb{R}$-valued process $L:=(L_{t})_{t\leq T}$ is a non-homogeneous Lévy process [^6] A non-homogeneous Lévy process $L$ can also be defined on a probability space without fixing a filtration a priori. In general, we work on a probability space equipped with the filtration generated by $L$; see [22, page 7] for details., if the following conditions hold:
  • $L$ has independent increments, i.e., $L_{t}-L_{s}$ is independent of $\mathcal{F}_{s}$ for any $0\leq s\leq t\leq T$.
  • For any $t\in[0,T]$, the law of $L_{t}$ is characterized by the characteristic function:
    $\mathbb{E}\left[e^{iuL_{t}}\right]=\exp\left\{\int_{0}^{t}\left(iub_{s}-\frac{c_{s}}{2}u^{2}+\int_{\mathbb{R}}\left(e^{iue}-1-iue\mathds{1}_{\{|e|\leq 1\}}\right)F_{s}(de)\right)ds\right\},$ (2.6)
    where $b:[0,T]\to\mathbb{R}$ and $c:[0,T]\to\mathbb{R}^{\ast,+}$ are Borel-measurable functions, and for each $s\in[0,T]$, $F_{s}$ is a measure on $\mathbb{R}$ that integrates $(1\wedge|e|^{2})$ and satisfies $F_{s}(\{0\})=0$. Moreover, for every $\mathsf{A}\in\mathcal{B}(\mathbb{R})$, the map $s\mapsto F_{s}(\mathsf{A})$ is Borel-measurable on $[0,T]$.
The triplet $(b,c,F):=(b_{t},c_{t},F_{t})_{t\leq T}$ is called the characteristics of $L$.
Remark 2.
  • In Definition 1, if the characteristics $(b_{t},c_{t},F_{t})$ are assumed to be independent of the time variable $t$, we obtain the classical definition of a homogeneous Lévy process.
  • From (2.6), we can easily derive that $L$ is a stochastically continuous process with $L_{0}=0$ a.s. Using this, along with the independent increments property, we conclude that every non-homogeneous Lévy process is an additive process.
In Definition 1, we did not impose regularity on the paths of $L$. However, since $L$ is an additive process (see Remark 2), it admits an RCLL modification (see [47, Ch 2. Theorem 11.5]). From now on, we consider this modification, ensuring that the left limits and jumps of $L$, denoted by $L_{t-}=\lim_{s\nearrow t}L_{s}$ and $\Delta L_{t}=L_{t}-L_{t-}$, are well-defined for all $t\in[0,T]$.
In order to obtain some necessary properties of the non-homogeneous Lévy process $L$, we will make the following assumptions for the remainder of the paper:
  • (NL1)
    The characteristics $(b_{t},c_{t},F)_{t\leq T}$ of $L$ satisfy:
    $\int_{0}^{T}\left(|b_{s}|+|c_{s}|+\int_{\mathbb{R}}\left(1\wedge|e|^{2}\right)F_{s}(de)\right)ds<+\infty$
  • (NL2)
    There are positive constants $\epsilon,\mathfrak{C}>0$ such that
    $\int_{0}^{T}\int_{\{|e|>1\}}e^{ue}F_{s}(de)ds<+\infty,\quad\forall u\in\left[-(1+\epsilon)\mathfrak{C},(1+\epsilon)\mathfrak{C}\right].$
The main needed properties of a non-homogeneous Lévy process under (NL1) and (NL2) are summarized in the following lemma:
Lemma 3.
  • Fix $t\in[0,T]$. Under assumption (NL1), the distribution of $L_{t}$ is infinitely divisible with the Lévy–Khintchine triplet $(\mathsf{b},\mathsf{c},\mathsf{F}):=(\mathsf{b}_{t},\mathsf{c}_{t},\mathsf{F}_{t})_{t\leq T}$ given by
    $\mathsf{b}_{t}=\int_{0}^{t}b_{s}\,ds,\quad\mathsf{c}_{t}=\int_{0}^{t}c_{s}\,ds,\quad\mathsf{F}_{t}(de)=\int_{0}^{t}F_{s}(de)\,ds.$
  • Under assumption (NL2), we have
    $\mathbb{E}\left[e^{uL_{t}}\right]<+\infty,\quad\quad\forall u\in\left[-(1+\epsilon)\mathfrak{C},(1+\epsilon)\mathfrak{C}\right].$
    As a consequence, the random variables $L_{t}$ have moments of all orders. In particular, the expected value of $L_{t}$ is finite.
  • Under assumptions (NL1) and (NL2), the process $L$ is a special $\mathbb{F}^{L}$-semimartingale with the canonical representation
    $L_{t}=\int_{0}^{t}\mathsf{b}_{s}\,ds+\int_{0}^{t}\sqrt{c_{s}}\,dW_{s}+\int_{0}^{t}\int_{\mathbb{R}}e\,\tilde{\mu}(ds,de).$ (2.7)
    Here,
    • $\mathsf{b}_{s}:=b_{s}+\int_{\mathbb{R}}e\mathds{1}_{\{\left|e\right|>1\}}F_{s}(de)$, with $b_{s}$ given in (2.6).
    • $\sqrt{c_{s}}$ is a measurable version of the square root of $c_{s}$.
    • $W:=(W_{t})_{t\leq T}$ is a standard one-dimensional Brownian motion.
    • $\tilde{\mu}:=\mu-\nu$ with $\mu$ is the random measure associated with the jumps of $L$ and $\nu$ is the compensator of $\mu$, defined by
      $\nu\left(\left[0,t\right]\times A\right):=\int_{0}^{t}\int_{A}F_{s}(de)\,ds,\quad A\in\mathcal{B}(\mathbb{R}).$ (2.8)
Note that the collection $(\mathcal{F}_{t})_{t\leq T}$ is neither increasing nor decreasing, so it does not constitute a filtration.
We define the power jumps of the non-homogeneous Lévy process $L$ by
$L_{t}^{(1)}=L_{t}\quad\mbox{and}\quad L_{t}^{(i)}=\sum_{0<s\leq t}(\Delta L_{s})^{i},\ i\geq 2.$
Let $m_{t}^{(1)}=\mathbb{E}[L_{t}]$ and $m_{t}^{(i)}=\mathbb{E}[L^{i}_{t}]=\int_{0}^{t}\int_{-\infty}^{+\infty}e^{i}\nu(de,ds)$ for $i\geq 2$ and $t\in[0,T]$, with $\nu$ defined by (2.8). Note that $m^{(1)}_{t}$ is well-defined due to Lemma 3.
Let us set $X_{t}^{(i)}=L_{t}^{(i)}-m_{t}^{(i)},\ i\geq 1$, the so-called Teugels martingales. For $i,j\geq 1$, the sharp bracket is given by $X^{(i)},X^{(j)}]_{t}=\int_{0}^{t}c_{s}ds\mathds{1}_{\{i=j=1\}}+L^{(i+j)}_{t}$ and the angle bracket is given by $\langle X^{(i)},X^{(j)}\rangle_{t}=\int_{0}^{t}c_{s}ds\mathds{1}_{\{i=j=1\}}+m^{(i+j)}_{t}$.
With the non-homogeneous Lévy process $(L_{t})_{t\leq T}$, we associate the family of processes $(H^{(i)})_{i\geq 1}$ defined by $H_{t}^{(i)}=\sum_{k=1}^{i}\alpha_{i,k}X_{t}^{(k)}$ where the coefficients $\alpha_{i,j}$ correspond to the orthonormalization of the polynomials $1,e,e^{2},\cdots$ with respect to the measure
$\pi([0,t],de)=\int_{0}^{t}c_{s}\delta_{0}(de)ds+\int_{0}^{t}e^{2}F_{s}(de)ds.$
More precisely, the polynomials $q_{n}$ defined by $q_{n}(e)=\sum_{k=1}^{n}\alpha_{n,k}e^{k-1}$ are orthonormal with respect to the measure $\pi$, i.e.
$\int_{0}^{t}\int_{-\infty}^{+\infty}q_{n}(e)q_{m}(e)\pi(ds,de)=0~\mbox{ if }n\neq m.$
We set
$p_{n}(e)=eq_{n}(e)=\alpha_{n,n}e^{n}+\alpha_{n,n-1}e^{n-1}+\cdots+\alpha_{n,1}e.$
The martingales $H^{(i)}$, called the orthonormalized $i$th-power-jump processes, are strongly orthogonal, meaning that $\langle H^{(i)},H^{(j)}\rangle=0$ for any $i\neq j$. Indeed, we have
$\langle H^{(i)},H^{(j)}\rangle_{t}$ $=$ $\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\alpha_{i,k}\alpha_{j,k^{\prime}}\left(\int_{0}^{t}\int_{-\infty}^{+\infty}e^{k+k^{\prime}}F_{s}(de)ds+\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=k^{\prime}=1\}}\right)$
$=$ $\int_{0}^{t}\int_{-\infty}^{+\infty}\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\alpha_{i,k}\alpha_{j,k^{\prime}}e^{k+k^{\prime}-2}\left\{e^{2}F_{s}(de)\right\}ds+\alpha_{i,1}\alpha_{j,1}\int_{0}^{t}c_{s}ds$
$=$ $\int_{0}^{t}\int_{-\infty}^{+\infty}q_{i}(e)q_{j}(e)\pi(de,ds).$
On the other hand, we have
$\begin{split}[H^{(i)},X^{(j)}]_{t}&=\sum_{k=1}^{i}\alpha_{i,k}\left(\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=j=1\}}+L^{(k+j)}_{t}\right)\\ &=\alpha_{i,i}L_{t}^{(i+j)}+\alpha_{i,i-1}L_{t}^{(i-1+j)}+\cdots+\alpha_{i,1}L_{t}^{(1+j)}+\alpha_{i,1}\int_{0}^{t}c_{s}ds\mathds{1}_{\{j=1\}}\end{split}$
Furthermore, for $i,j\geq 1$
$H^{(i)},H^{(j)}]_{t}=\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\left(\alpha_{i,k}\alpha_{j,k^{\prime}}L_{t}^{(k+k^{\prime})}+\alpha_{i,1}^{2}\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=k^{\prime}=1\}}\right).$
Throughout this paper, we restrict our analysis to the case where the non-homogeneous Lévy process $L$ admits a finite number, d1d-1, of jump sizes for some $d\geq 1$. We set (see, e.g., [8, Section 2.3, Example 3])
$d\langle H^{(k)},H^{(k^{\prime})}\rangle_{t}=|\gamma^{(k)}_{t}||\gamma^{(k^{\prime})}_{t}|\mathsf{D}_{k,k^{\prime}}dt\quad\text{for all}\quad t\in[0,T]\quad\text{and}\quad{k,k^{\prime}\in\{1,\dots,d\}},$
where $(\gamma_{t})_{t\leq T}=\left(\gamma^{(1)}_{t},\gamma^{(2)}_{t},\cdots,\gamma^{(d)}_{t}\right)_{t\leq T}$ are deterministic continuous functions and $\mathsf{D}_{k,j}$ is the Kronecker delta function of $k$ and $j$.
3 RGBDSDE-NL with stochastic monotone coefficients
This section presents the first main result of the paper. We begin by stating the assumptions on the data $(\xi,f,h,\kappa,g,S)$ of the RGBDSDE-NL (1.5). We then provide an auxiliary result: an extension of the classical Itô formula, with a full description and proof. This extension yields key a priori estimates that are used throughout the paper to establish uniqueness and to streamline several calculations. Finally, under stochastic monotonicity and Lipschitz conditions on the drivers, we establish existence and uniqueness via Yosida regularization and Picard iteration.
3.1 Assumptions and technical results
Let us now state the assumption imposed on the data of the RGBDSDE-NL (1.5).
Assumptions on the data $(\xi,f,h,\kappa,g,S)$
First, we define the space $\ell^{2}=\big\{\mathsf{z}=\{\mathsf{z}_{k}\}_{1\leq k\leq d}\in\mathbb{R}^{d};\ \|\mathsf{z}\|_{\ell^{2}}^{2}=\sum_{k=1}^{d}|\mathsf{z}_{k}|^{2}<+\infty\big\}$.
- Measurability of the data and trajectory properties of the process $(\kappa_{t})_{t\leq T}$:
  • $\xi$ is an $\mathcal{F}_{T}$-measurable random variable.
  • The process $(\kappa_{t})_{t\leq T}$ is a continuous non-decreasing process such that $\kappa_{0}=0$ and $\kappa_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$;
  • $f$, $h$ and $g$ are jointly measurable;
  • $\forall t\in[0,T]$, $y\in\mathbb{R}$, $z\in\ell^{2}$, the processes $f\left(\cdot,t,y,z\right):\Omega\rightarrow\mathbb{R}$, $g\left(\cdot,t,y,z\right):\Omega\rightarrow\mathbb{R}$, and $g\left(\cdot,t,y\right):\Omega\rightarrow\mathbb{R}$ are $\mathcal{F}_{t}$-measurable.
- Stochastic Monotonicity of $f$ and $h$ in $y$:
There exist two processes $\lambda:\Omega\times[0,T]\rightarrow\mathbb{R}$ and $\varrho:\Omega\times[0,T]\rightarrow\mathbb{R}^{\ast}_{-}$ such that
  • (i)
    $\lambda_{t}$ and $\varrho_{t}$ are $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
  • (ii)
    For all $y$, $y^{\prime}\in\mathbb{R}$, $z\in\ell^{2}$, $d\mathbb{P}\otimes dt$-a.e.,
    $\left(y-y^{\prime}\right)\left(f\left(t,y,z\right)-f\left(t,y^{\prime},z\right)\right)\leq\lambda_{t}\left|y-y^{\prime}\right|^{2}.$
  • (iii)
    For all $y$, $y^{\prime}\in\mathbb{R}$, $d\mathbb{P}\otimes d\kappa_{t}$-a.e.,
    $\left(y-y^{\prime}\right)\left(h\left(t,y\right)-h\left(t,y^{\prime}\right)\right)\leq\varrho_{t}\left|y-y^{\prime}\right|^{2}.$
- Stochastic Lipschitz condition on $f$ in $z$:
There exists a stochastic process $\eta:\Omega\times[0,T]\rightarrow\mathbb{R}^{\ast}_{+}$ such that
  • (iv)
    $\eta_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$,
  • (v)
    For all $y\in\mathbb{R}$, $z$, $z^{\prime}\in\ell^{2}$, $d\mathbb{P}\otimes dt$-a.e.,
    $\left|f\left(t,y,z\right)-f\left(t,y,z^{\prime}\right)\right|\leq\eta_{t}\left\|\gamma_{t}\left(z-z^{\prime}\right)\right\|_{\ell^{2}}.$
- Stochastic Lipschitz condition on $g$ in $y$ and a Lipschitz condition in $z$:
There exists a stochastic process $\rho:\Omega\times[0,T]\rightarrow\mathbb{R}^{\ast}_{+}$ and a constant $\alpha\in(0,\frac{1}{2})$ such that
  • (vi)
    $\rho_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
  • (vii)
    For all $y$, $y^{\prime}\in\mathbb{R}$, $z$, $z^{\prime}\in\ell^{2}$, $d\mathbb{P}\otimes dt$-a.e.,
    $\left|g\left(t,y,z\right)-g\left(t,y^{\prime},z^{\prime}\right)\right|^{2}\leq\rho_{t}\left|y-y^{\prime}\right|^{2}+\alpha\left\|\gamma_{t}\left(z-z^{\prime}\right)\right\|^{2}_{\ell^{2}}.$
- Linear growth of $f$ and $h$:
For some constant $\zeta>0$, a $\left[0,\infty\right)$-valued process $(\phi_{t})_{t\leq T}$, and some $\left[1,\infty\right)$-valued processes $\left\{\varphi_{t},\psi_{t};0\leq t\leq T\right\}$, such that for all $(t,y)\in[0,T]\times\mathbb{R}$, we have
  • (viii)
    $\varphi_{t}$, $\phi_{t}$ and $\psi_{t}$ are $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$;
  • (ix)
    $\left|f(t,y,0)\right|\leq\varphi_{t}+\phi_{t}\left|y\right|$ and $\left|h(t,y)\right|\leq\psi_{t}+\zeta\left|y\right|$.
- Continuity condition on $f$ and $g$:
For all $z\in\ell^{2}$:
  • (x)
    $d\mathbb{P}\otimes dt$-a.e., the mapping $y\mapsto f\left(t,y,z\right):\mathbb{R}\rightarrow\mathbb{R}$ is continuous.
  • (xi)
    $d\mathbb{P}\otimes d\kappa_{t}$-a.e., the mapping $y\mapsto h\left(t,y\right):\mathbb{R}\rightarrow\mathbb{R}$ is continuous.
-Obstacle and integrability assumption:
We consider a stochastic continuous process $S:=(S_{t})_{t\leq T}$ satisfying
  • $S_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
  • $\xi\leq S_{T}$.
    Throughout the current paper, the process $S$ will be referred to as a barrier or an obstacle.
Let $(V_{t})_{t\leq T}$ and $(Q_{t})_{t\leq T}$ be the two continuous non-deceasing stochastic processes defined by
$V_{t}:=\int_{0}^{t}a_{s}^{2}ds,\quad\mathcal{Q}_{t}:=V_{t}+\kappa_{t},$
with $a^{2}_{s}:=\left|\lambda_{s}\right|+\phi_{s}+\phi^{2}_{s}+\rho_{s}+\eta^{2}_{s}$. We assume that, there exists some $\epsilon>0$, such that $a^{2}_{s}\geq\epsilon$. Since $\varrho_{s}<0$, we can also consider $\left|\varrho_{s}\right|\geq\epsilon$.
Let us set
$\phi^{\theta,\mu}_{t}=\theta V_{t}+\mu\kappa_{t}\mbox{ and }\Phi^{\theta,\mu}:=e^{\phi^{\theta,\mu}}\mbox{ for }\theta,\mu>0.$
We assume that, for any $\theta$, $\mu>0$
  • (xii)
    $\mathbb{E}\left[\Phi^{\theta,\mu}_{T}|\xi|^{2}\right]<+\infty.$
  • (xiii)
    $\mathbb{E}\left[\int_{0}^{T}{\Phi^{\theta,\mu}_{t}}\left(\left|\frac{\varphi_{t}}{a_{t}}\right|^{2}+\left|g(t,0,0)\right|^{2}\right)dt\right]+\mathbb{E}\left[\int_{0}^{T}{\Phi^{\theta,\mu}_{t}}\psi_{t}^{2}d\kappa_{t}\right]<+\infty$.
  • (xiv)
    $\mathbb{E}\left[\sup_{t\in[0,T]}\big|\Phi^{\theta,\mu}_{t}\left(S^{+}_{t}\right)\big|^{2}\right]<+\infty$.
In the rest of this paper, the previous assumptions will be denoted by (H-M).
For any $\theta,\mu>0$, we consider the following spaces:
  • $\mathcal{S}^{2}_{\theta,\mu}$: The space of RCLL processes $Y:\Omega\times[0,T]\longrightarrow\mathbb{R}$ such that
    • $Y_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
    • ${\left\|Y\right\|^{2}_{\mathcal{S}^{2}_{\theta,\mu}}=\mathbb{E}\left[\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}|Y_{t}|^{2}\right]<+\infty.}$
      (with the convention $\mathcal{S}^{2}_{\mu}:=\mathcal{S}^{2}_{0,\mu}$ and $\mathcal{S}^{2}:=\mathcal{S}^{2}_{0,0}$).
  • $\mathcal{A}^{2}$: The space of continuous non-decreasing processes $K:\Omega\times[0,T]\longrightarrow\mathbb{R}$ such that $K_{0}=0$ and
    • $K_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
    • ${\left\|K\right\|^{2}_{\mathcal{A}^{2}}=\mathbb{E}\left|K_{T}\right|^{2}<+\infty.}$
  • $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}$: The space of jointly measurable stochastic processes $Y:\Omega\times[0,T]\longrightarrow\mathbb{R}$ such that
    • $Y_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
    • ${\left\|Y\right\|^{2}_{\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}}=\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{t}|Y_{t}|^{2}dQ_{t}<+\infty.}$
      (with the convention $\mathcal{H}^{2,\mathcal{Q}}_{\mu}:=\mathcal{H}^{2,\mathcal{Q}}_{0,\mu}$, $\mathcal{H}^{2,\mathcal{Q}}:=\mathcal{H}^{2,\mathcal{Q}}_{0,0}$, and $\mathcal{H}^{2,\kappa}_{\mu}$ is the same space as $\mathcal{H}^{2,\mathcal{Q}}_{0,\mu}$, with integrability taken with respect to $\kappa$).
  • $\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$: The space of jointly measurable $\ell^{2}$-valued processes $(Z_{t})_{t\leq T}=\left(\{Z^{k}_{t}\}_{1\leq k\leq d}\right)_{t\leq T}$ such that
    • $Z$ is $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$.
    • ${\|Z\|^{2}_{\mathcal{H}^{2,\ell^{2}}_{\beta}}=\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{t}\|\gamma_{t}Z_{t}\|_{\ell^{2}}^{2}dt=\sum_{k=1}^{d}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{t}|Z_{t}^{(k)}|^{2}|\gamma_{t}^{(k)}|^{2}dt<+\infty.}$
      (with the convention $\mathcal{H}^{2,\ell^{2}}_{\mu}:=\mathcal{H}^{2,\ell^{2}}_{\mu,0}$ and $\mathcal{H}^{2,\ell^{2}}:=\mathcal{H}^{2,\ell^{2}}_{0,0}$).
Notice that the space $\mathcal{C}^{2}_{\theta,\mu}=\mathcal{S}^{2}_{\theta,\mu}\cap\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}$ endowed with the norm $\left\|Y\right\|^{2}_{\mathcal{C}^{2}_{\theta,\mu}}=\left\|Y\right\|_{\mathcal{S}^{2}_{\theta,\mu}}^{2}+\left\|Y\right\|_{\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}}^{2}$ is a Banach space as is the space $\mathcal{B}^{2}_{\theta,\mu}=\mathcal{C}^{2}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}\times\mathcal{A}^{2}$ with the norm $\left\|(Y,Z,K)\right\|^{2}_{\mathcal{B}^{2}_{\theta,\mu}}=\left\|Y\right\|_{\mathcal{S}^{2}_{\theta,\mu}}^{2}+\left\|Y\right\|_{\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}}^{2}+\left\|Z\right\|_{\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}}^{2}+\left\|K\right\|_{\mathcal{A}^{2}}^{2}.$
Definition 4.
Let $\theta,\mu>0$. A solution to the RGBDSDE-NL (1.5) associated with the parameters $(\xi,f,h,\kappa,g,S)$ is a triplet of processes $(Y,Z,K)$ that satisfies (1.5) and belongs to the space $\mathcal{B}^{2}_{\theta,\mu}$.
Remark 5.
Note that, since $\left\{\varphi_{t},\psi_{t};\ 0\leq t\leq T\right\}$ are $1,\infty)-$valued processes, then
$\mathbb{E}\int_{0}^{T}{\Phi^{\theta,\mu}_{s}}dQ_{s}<\infty.$
In the following, let $\mathfrak{c}>0$ denotes a finite constant that may take different values from line to line. Additionally, we will use the notation $\mathfrak{c}_{\nu}$ to indicate that the constant $\mathfrak{c}$ depends on a specific set of parameters $\nu$. Finally, it should be noted that references to assumptions (H-M)-(i)-(ix)-(x) are equivalent to referencing (H-M)-(i), (H-M)-(ix) and (H-M)-(x) separately.
The following extension of the well-known Itô formula is crucial for the remainder of the paper, particularly in deriving useful a priori estimates between the solutions of the RGBDSDE-NL (1.5). The proof is analogous to the one provided in [36, Lemma 1.3], and is based on the techniques outlined in [42, Ch. II, Theorem 32].
Lemma 6.
Let $\mathcal{Y}\in\mathcal{S}^{2}$, $\mathsf{B},\mathsf{C},\mathsf{G}\in\mathcal{H}^{2,\mathcal{Q}}$, $\mathcal{K}\in\mathcal{A}^{2}$, and $\mathsf{Z}\in\mathcal{H}^{2,\ell}$ be such that
$\mathcal{Y}_{t}=\mathcal{Y}_{0}+\int_{0}^{t}\mathsf{B}_{s}\,ds+\int_{0}^{t}\mathsf{C}_{s}\,d\kappa_{s}+\int_{0}^{t}\mathsf{G}_{s}\,\overleftarrow{dB}_{s}+\mathcal{K}_{t}+\sum_{k=1}^{d}\int_{0}^{t}\mathsf{Z}_{s}^{(k)}\,dH_{s}^{(k)}.$ (3.9)
Then, for any $\Psi\in\mathcal{C}^{1,1,2}\left(\mathbb{R}^{3}\right)$ and any 2-dimensional continuous non-decreasing process $\left(\mathsf{A}_{t}\right)_{t\leq T}=\left(\mathsf{V}_{t},\mathsf{K}_{t}\right)_{t\leq T}$ such that $\mathsf{V}_{0}=\mathsf{K}_{0}=0$ and $\mathsf{V}_{t},\mathsf{K}_{t}\in\mathcal{F}_{t}$ for all $t\in[0,T]$, we have
$\Psi\left(\mathsf{A}_{t},\mathcal{Y}_{t}\right)$ $=\Psi\left(0,\mathcal{Y}_{0}\right)+\int_{0}^{t}\frac{\partial\Psi}{\partial x}\left(\mathsf{A}_{s},\mathcal{Y}_{s}\right)d\mathsf{V}_{s}+\int_{0}^{t}\frac{\partial\Psi}{\partial y}\left(\mathsf{A}_{s},\mathcal{Y}_{s}\right)d\mathsf{K}_{s}+\int_{0}^{t}\frac{\partial\Psi}{\partial z}\left(\mathsf{A}_{s},\mathcal{Y}_{s}\right)\mathsf{B}_{s}\,ds$
$+\int_{0}^{t}\frac{\partial\Psi}{\partial z}\left(\mathsf{A}_{s},\mathcal{Y}_{s}\right)\mathsf{C}_{s}\,d\kappa_{s}+\int_{0}^{t}\frac{\partial\Psi}{\partial z}\left(\mathsf{A}_{s},\mathcal{Y}_{s}\right)\mathsf{G}_{s}\,\overleftarrow{dB}_{s}+\int_{0}^{t}\frac{\partial\Psi}{\partial z}\left(\mathsf{A}_{s},\mathcal{Y}_{s}\right)\,d\mathcal{K}_{s}$
$+\sum_{k=1}^{d}\int_{0}^{t}\frac{\partial\Psi}{\partial z}\left(\mathsf{A}_{s},\mathcal{Y}_{s-}\right)\mathsf{Z}_{s}^{(k)}\,dH_{s}^{(k)}-\frac{1}{2}\int_{0}^{t}\frac{\partial^{2}\Psi}{\partial z^{2}}\left(\mathsf{A}_{s},\mathcal{Y}_{s-}\right)\left|\mathsf{G}_{s}\right|^{2}\,ds$
$+\frac{1}{2}\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\frac{\partial^{2}\Psi}{\partial z^{2}}\left(\mathsf{A}_{s},\mathcal{Y}_{s-}\right)\mathsf{Z}_{s}^{(k)}\mathsf{Z}_{s}^{(k^{\prime})}\,d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}$
$+\sum_{0<s\leq t}\left\{\Psi(\mathsf{A}_{s},\mathcal{Y}_{s})-\Psi(\mathsf{A}_{s},\mathcal{Y}_{s-})-\frac{\partial\Psi}{\partial z}\left(\mathsf{A}_{s},\mathcal{Y}_{s-}\right)\Delta\mathcal{Y}_{s}-\frac{1}{2}\frac{\partial^{2}\Psi}{\partial z^{2}}\left(\mathsf{A}_{s},\mathcal{Y}_{s-}\right)\left(\Delta\mathcal{Y}_{s}\right)^{2}\right\},$
with $\Delta\mathcal{Y}_{s}=\sum_{k=1}^{d}\mathsf{Z}_{s}^{(k)}\,\Delta H_{s}^{(k)}$.
Corollary 7.
Let $\mathcal{Y}$ be a stochastic process of the form (3.9), and let $\theta,\mu>0$. Then, for every $t\in[0,T]$, almost surely, we have
$\begin{split}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}&=\left|\mathcal{Y}_{0}\right|^{2}+\theta\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{0}\right|^{2}dV_{s}+\mu\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{0}\right|^{2}d\kappa_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{B}_{s}ds\\ &+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{C}_{s}d\kappa_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{G}_{s}\overleftarrow{dB}_{s}+2\sum_{k=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s-}\mathsf{Z}_{s}^{(k)}dH_{s}^{(k)}\\ &-\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds+\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\mathsf{Z}_{s}^{(k)}\mathsf{Z}_{s}^{(k^{\prime})}\,d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}d\mathcal{K}_{s}.\end{split}$
Moreover, if $\left(\mathcal{Y},\mathsf{Z},\mathsf{B},\mathsf{C},\mathsf{G}\right)\in\left(\mathcal{S}^{2}_{\theta,\mu}\cap\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\right)\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}\times\left(\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\right)^{3}$, then using $d\langle H^{(k)},H^{(k^{\prime})}\rangle_{t}=|\gamma^{(k)}_{t}||\gamma^{(k^{\prime})}_{t}|\mathsf{D}_{k,k^{\prime}}\,dt$, we have
$\begin{split}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}&=\mathbb{E}\left|\mathcal{Y}_{0}\right|^{2}+\theta\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{s}\right|^{2}d\kappa_{s}+2\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{B}_{s}ds\\ &+2\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{C}_{s}d\kappa_{s}-\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds+\sum_{k=1}^{d}\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{Z}_{s}^{(k)}\gamma^{(k)}_{s}\right|^{2}ds+\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}d\mathcal{K}_{s}.\end{split}$
Proof.
For the first part, it suffices to apply Lemma 6 with $\Psi(x,y,z)=e^{\theta x+\mu y}\left|z\right|^{2}$ and $\left(\mathsf{A}_{t}\right)_{t\leq T}=(V_{t},\kappa_{t})_{t\leq T}$. Following this, in the expectation calculation stage, we apply the Burkholder-Davis-Gundy inequality (see, e.g., [42, Ch IV. Theorem 48]), which implies
$\begin{split}\mathbb{E}\sup_{0\leq t\leq T}\left|\int_{t}^{T}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{G}_{s}\overleftarrow{dB}_{s}\right|&\leq\mathfrak{c}\mathbb{E}\sqrt{\int_{0}^{T}\Phi^{2\theta,2\mu}_{s}\mathcal{Y}_{s}^{2}\mathsf{G}_{s}^{2}ds}\\ &\leq\frac{\mathfrak{c}}{2}\left(\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds\right)<+\infty.\end{split}$
Hence, $\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{G}_{s}\overleftarrow{dB}_{s}=0$ for any $t\in[0,T]$. Similarly, we can show that $\mathbb{E}\sum_{k=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s-}\mathsf{Z}_{s}^{(k)}dH_{s}^{(k)}=0$ for any $t\in[0,T]$, using the following:
$\begin{split}\mathbb{E}\sup_{0\leq t\leq T}\left|\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s-}\mathsf{Z}^{(k)}_{s}dH^{(k)}_{s}\right|&\leq\mathfrak{c}\mathbb{E}\sqrt{\sum_{k=1}^{d}\int_{0}^{T}\Phi^{2\theta,2\mu}_{s}\mathcal{Y}_{s}^{2}\left|\mathsf{Z}^{(k)}_{s}\right|^{2}\left|\gamma^{(k)}_{s}\right|^{2}ds}\\ &\leq\frac{\mathfrak{c}}{2}\left(\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}+\mathbb{E}\sum_{k=1}^{d}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathsf{Z}^{(k)}_{s}\right|^{2}\left|\gamma^{(k)}_{s}\right|^{2}ds\right)<+\infty.\end{split}$
3.2 A priori estimates and uniqueness
In this part, we provide some a priori estimates for the solutions of the RGBDSDE-NL (1.5), which will be useful for further applications.
Proposition 8.
Assume that (H-M) holds. Let $\theta,\mu>0$. Let $(Y,Z,K)$ be a solution to the RGBDSDE-NL (1.5) with parameters $(\xi,f,h,g,\kappa,S)$. Then, there exists a constant $\theta_{0}>4+\frac{2}{1-2\alpha}$ such that for any $\theta>\theta_{0}$ and $\mu>0$, we have
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\mathbb{E}\left[\left|K_{T}\right|^{2}\right]\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}\left(S_{t}\right)^{+}\right|^{2}\right).\end{split}$
Proof.
Using Corollary 7, we have
$\begin{split}&\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\theta\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mu\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &=\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}f(s,Y_{s},Z_{s})ds+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}h(s,Y_{s})d\kappa_{s}\\ &\quad+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds-2\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s-}Z^{(k)}_{s}dH^{(k)}_{s}\\ &\quad-\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}Z_{s}^{(k)}Z_{s}^{(k^{\prime})}\,\left(d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}-d\big\langle H^{(k)},H^{(k^{\prime})}\big\rangle_{s}\right)+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}dK_{s}.\end{split}$ (3.10)
Using (H-M)-(ii)-(iii)-(v)-(vii) and the basic inequality $2ab\leq\varepsilon a^{2}+\frac{1}{\varepsilon}b^{2}$ for any $\varepsilon>0$, we derive
$\begin{split}2Y_{s}f(s,Y_{s},Z_{s})ds&\leq 2\lambda_{s}\left|Y_{s}\right|^{2}ds+2\left|Y_{s}\right|\left(\varphi_{s}+\eta_{s}\left\|\gamma_{s}Z_{s}\right\|_{\ell^{2}}\right)ds\\ &\leq\left\{2\lambda_{s}+\frac{1}{\varepsilon^{\prime}}\eta_{s}^{2}+\varepsilon a^{2}_{s}\right\}\left|Y_{s}\right|^{2}ds+\frac{1}{\varepsilon}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\varepsilon^{\prime}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\left(2+\varepsilon+\frac{1}{\varepsilon^{\prime}}\right)\left|Y_{s}\right|^{2}dV_{s}+\frac{1}{\varepsilon}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\varepsilon^{\prime}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds,\end{split}$
and
$\begin{split}2Y_{s}h(s,Y_{s})d\kappa_{s}&\leq 2\varrho_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+2\left|Y_{s}\right|\psi_{s}d\kappa_{s}\\ &\leq\varepsilon^{\prime\prime}\left|Y_{s}\right|^{2}d\kappa_{s}+\frac{1}{\varepsilon^{\prime\prime}}\varphi_{s}^{2}d\kappa_{s},\end{split}$
and
$\begin{split}\left|g(s,Y_{s},Z_{s})\right|^{2}ds&\leq 2\left(\rho_{s}\left|Y_{s}\right|^{2}+\alpha\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\right)+2\left|g(s,0,0)\right|^{2}ds\\ &\leq 2\left|Y_{s}\right|^{2}dV_{s}+2\alpha\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds+2\left|g(s,0,0)\right|^{2}ds.\end{split}$
Then, plugging those in (3.10), using the Skorokhod condition and taking expectation, we get
$\begin{split}&\mathbb{E}\left[\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}\right]+\left(\theta-4-\varepsilon-\frac{1}{\varepsilon^{\prime}}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}\\ &\qquad+\left(\mu-\varepsilon^{\prime\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+(1-2\alpha-\varepsilon^{\prime})\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathbb{E}\left[\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}\right]+\frac{1}{\varepsilon}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\frac{1}{\varepsilon^{\prime\prime}}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\\ &\qquad+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}S_{s}dK_{s}.\end{split}$
Set $\theta_{0}=4+\epsilon+\frac{1}{\varepsilon^{\prime}}$. Choosing $\varepsilon=\frac{1}{1-2\alpha}$, $\theta>\theta_{0}$ , $\varepsilon^{\prime}<1-2\alpha$, and $\varepsilon^{\prime\prime}=\frac{\mu}{2}$, one has for any $\varepsilon^{\prime\prime\prime}>0$
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\right)\\ &\qquad+\varepsilon^{\prime\prime\prime}\mathfrak{c}_{\theta,\mu,\alpha}^{2}\mathbb{E}\sup_{t\leq u\leq T}\left|\Phi^{\theta,\mu}_{u}(S_{u})^{+}\right|^{2}+\frac{1}{\varepsilon^{\prime\prime\prime}}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}.\end{split}$ (3.11)
Coming back to (1.5)-(i), by squaring, we have
$\begin{split}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}&\leq 6\left(\mathbb{E}\left|Y_{t}\right|^{2}+\mathbb{E}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}\right.\\ &\left.+\mathbb{E}\left(\int_{t}^{t}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}\right)^{2}+\mathbb{E}\left(\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s}\right)^{2}\right)\\ &\leq 6\left(\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}\right.\\ &\qquad\left.+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right).\end{split}$ (3.12)
Using (H-M)-(vii)-(ix) and Hölder’s inequality we have
$\begin{split}\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}&\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{-\theta,\mu}_{s}dV_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{f(s,Y_{s},Z_{s})}{a_{s}}\right|^{2}ds\right)\\ &\leq\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\frac{\varphi^{2}_{s}+\phi_{s}^{2}\left|Y_{s}\right|^{2}+\eta_{s}^{2}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}}{a^{2}_{s}}ds\\ &\leq\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\end{split}$
and
$\begin{split}\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}&\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{\theta,-\mu}_{s}d\kappa_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|h(s,Y_{s})\right|^{2}d\kappa_{s}\right)\\ &\leq\frac{2\zeta^{2}}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+\frac{2}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\psi^{2}_{s}d\kappa_{s}.\end{split}$
Next, using ,we have
$\begin{split}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds&\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(\left|Y_{s}\right|^{2}dV_{s}+\alpha\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right)+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds.\end{split}$
Therefore
$\begin{split}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}&\leq 6\mathbb{E}\left|Y_{t}\right|^{2}+6\mathbb{E}\left|\xi\right|^{2}+\left(\frac{18}{\theta}+\frac{12\zeta^{2}}{\mu}+12\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}\\ &\qquad+\left(\frac{18}{\theta}+6\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\frac{18}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds\\ &\qquad+12\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds.\end{split}$ (3.13)
Plugging this into (3.11), choosing $\varepsilon^{\prime\prime\prime}>\frac{18}{\theta}+\frac{12\zeta^{2}}{\mu}+12$, we have
$\begin{split}&\sup_{0\leq t\leq T}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$ (3.14)
Then, from (3.13), we deduce
$\begin{split}\mathbb{E}\left|K_{T}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$
Now, applying the Burkholder-Davis–Gundy inequality similar to those used in Corollary 7, we have
$\begin{split}2\mathbb{E}\sup_{0\leq t\leq T}\left|\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}\right|&\leq 2\mathfrak{c}\mathbb{E}\sqrt{\int_{0}^{T}\Phi^{2\theta,2\mu}_{s}Y_{s}^{2}\left|g(s,Y_{s},Z_{s})\right|^{2}ds}\\ &\leq\frac{1}{4}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}+4\mathfrak{c}^{2}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds\\ &\leq\frac{1}{4}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}+8\mathfrak{c}^{2}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mathfrak{c}^{2}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\end{split}$
and
$\begin{split}2\mathbb{E}\sup_{0\leq t\leq T}\left|\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s-}Z^{(k)}_{s}dH^{(k)}_{s}\right|&\leq 2\mathfrak{c}\mathbb{E}\sqrt{\sum_{k=1}^{d}\int_{0}^{T}\Phi^{2\theta,2\mu}_{s}Y_{s}^{2}\left|Z^{(k)}_{s}\right|^{2}\left|\gamma^{(k)}_{s}\right|^{2}ds}\\ &\leq\frac{1}{4}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+4\mathfrak{c}^{2}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds.\end{split}$
Finally, the Burkholder-Davis-Gundy inequality along with the Kunita-Watanabe inequality (see, e.g., [42, Ch II. Theorem 25]), we get
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\left|\sum_{k,k^{\prime}=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}Z^{(k)}_{s}Z^{(k^{\prime})}_{s}\left(d[H^{(k)},H^{(k^{\prime})}]_{s}-d\langle H^{(k)},H^{(k^{\prime})}\rangle_{s}\right)\right|\\ &\leq\mathfrak{c}\mathbb{E}\sqrt{\sum_{k,k^{\prime}=1}^{d}\sum_{0<s\leq T}\Phi^{\theta,\mu}_{s}\left|Z^{(k)}_{s}\right|^{2}\left|Z^{(k^{\prime})}_{s}\right|^{2}\left(\Delta H^{(k)}_{s}\right)^{2}\left(\Delta H^{(k^{\prime})}_{s}\right)^{2}}.\\ &\leq\mathfrak{c}\mathbb{E}\sum_{k,k^{\prime}=1}^{d}\sum_{0<s\leq T}\Phi^{\theta,\mu}_{s}\left|Z^{(k)}_{s}\right|\left|Z^{(k^{\prime})}_{s}\right|\left|\Delta H^{(k)}_{s}\right|\left|\Delta H^{(k^{\prime})}_{s}\right|.\\ &\leq\mathfrak{c}\mathbb{E}\sum_{k,k^{\prime}=1}^{d}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Z^{(k)}_{s}\right|\left|Z^{(k^{\prime})}_{s}\right|d\left\|\left[H^{(k)},H^{(k^{\prime})}\right]\right\|_{s}.\\ &\leq\mathfrak{c}\mathbb{E}\sum_{k,k^{\prime}=1}^{d}\left(\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Z^{(k)}_{s}\right|^{2}d\left[H^{(k)},H^{(k)}\right]_{s}\right)^{\frac{1}{2}}\left(\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Z^{(k^{\prime})}_{s}\right|^{2}d\left[H^{(k^{\prime})},H^{(k^{\prime})}\right]_{s}\right)^{\frac{1}{2}}.\\ &\leq\mathfrak{c}\mathbb{E}\sum_{k=1}^{d}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Z^{(k)}_{s}\right|^{2}d\left[H^{(k)},H^{(k)}\right]_{s}+\mathfrak{c}\mathbb{E}\sum_{k^{\prime}=1}^{d}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Z^{(k^{\prime})}_{s}\right|^{2}d\left[H^{(k^{\prime})},H^{(k^{\prime})}\right]_{s}.\\ &\leq\mathfrak{c}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\|\gamma_{s}Z_{s}\|_{\ell^{2}}^{2}ds,\end{split}$
where $\left\|\left[H^{(k)},H^{(k^{\prime})}\right]\right\|_{s}$ is the total variation of $\left[H^{(k)},H^{(k^{\prime})}\right]_{s}$.
Returning to (3.10), taking supremum and using the above three inequalities, along with (3.14), we derive the following
$\begin{split}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$
Completing the proof. ∎
Proposition 9.
Assume that (H-M) hold. Then, the RGBDSDE-NL (1.5) has at most one solution $(Y,Z,K)$.
Proof.
Let $(Y,Z,K)$ and $(Y^{\prime},Z^{\prime},K^{\prime})$ be two solutions of the RGBDSDE-NL (1.5) associated with the data $(\xi,f,h,\kappa,g,S)$. Denote $\bar{\mathfrak{S}}=\mathfrak{S}-\mathfrak{S}^{\prime}$ for $\mathfrak{S}\in\left\{Y,Z,K\right\}$.
Using Corollary 7 and the minimality condition $\bar{Y}_{s}d\bar{K}_{s}\leq 0$, we obtain the following result for any $\theta,\mu>0$:
$\begin{split}&\Phi^{\theta,\mu}_{t}\left|\bar{Y}_{t}\right|^{2}+\theta\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\mu\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}d\kappa_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq 2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}f(s,Y_{s},Z_{s})ds+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}\left(h(s,Y_{s})-h(s,Y^{\prime}_{s})\right)d\kappa_{s}\\ &\quad+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(g(s,Y_{s},Z_{s})-g(s,Y^{\prime}_{s},Z^{\prime}_{s})\right)\overleftarrow{dB}_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})-g\left(s,Y^{\prime}_{s},Z^{\prime}_{s}\right)\right|^{2}ds\\ &\quad-2\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s-}\bar{Z}^{(k)}_{s}dH^{(k)}_{s}-\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\bar{Z}_{s}^{(k)}\bar{Z}_{s}^{(k^{\prime})}\,\left(d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}-d\big\langle H^{(k)},H^{(k^{\prime})}\big\rangle_{s}\right).\end{split}$
Then, by using (H-M)-(ii)-(iii)-(v)-(vii) and taking the expectation, we obtain:
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\bar{Y}_{t}\right|^{2}+\theta\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(f(s,Y_{s},Z_{s})-f(s,Y^{\prime}_{s},Z^{\prime}_{s})\right)ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\bar{Y}_{s}\left(h(s,Y_{s})-h(s,Y^{\prime}_{s})\right)d\kappa_{s}\\ &\quad+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})-g(s,Y^{\prime}_{s},Z^{\prime}_{s})\right|^{2}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\left(\alpha+\frac{1}{2}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds.\end{split}$
Then, for $\theta>2$ and $\mu>0$, and using the fact that $\alpha+\frac{1}{2}<1$, we have:
$\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\bar{Y}_{t}\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}dV_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\bar{Y}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|\bar{Z}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\leq 0.$ (3.15)
Therefore, we have $Z=Z^{\prime}$. On the other hand, by applying the Burkholder-Davis-Gundy inequality and performing similar calculations as those in Proposition 8, together with (3.15), we derive that $\left\|\bar{Y}\right\|^{2}_{\mathcal{S}^{2}_{\theta,\mu}}=0$ for any $(\theta,\mu)\in(2,+\infty)\times(0,+\infty)$. Thus, $Y=Y^{\prime}$ (in the indistinguishable sense). Finally, from (1.5)-(i), we have
$K_{t}=Y_{0}-Y_{t}-\int_{0}^{t}f(s,Y_{s},Z_{s})\,ds-\int_{0}^{t}h(s,Y_{s})\,d\kappa_{s}-\int_{0}^{t}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}+\sum_{k=1}^{d}\int_{0}^{t}Z^{(k)}_{s}\,dH^{(k)}_{s}$.
Hence, $K=K^{\prime}$, completing the proof. ∎
3.3 Existence
The proof of existence will be divided into two parts:
  • In the first part, we deal with a special case where the generator $f$ does not depend on the variable $z$ and $g$ does not depend on the variables $(y,z)$, using Yosida approximation of monotone functions.
  • In the second part, which addresses the general case, we resolve the problem using the Picard iteration method.
3.3.1 Case where generators $f$ do not depend on $z$, and $g$ does not depend on $(y,z)$
We consider the following version of the RGBDSDE-NL (1.5):
$\left\{\begin{split}\text{(i)}&~Y_{t}=\xi+\int_{t}^{T}\mathfrak{f}(s,Y_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}+\int_{t}^{T}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\left(K_{T}-K_{t}\right)-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s},\\ \text{(ii)}&~Y_{t}\geq S_{t},~t\in[0,T]~\mbox{ and }~\int_{0}^{T}\left(Y_{s}-S_{s}\right)dK_{s}=0.\end{split}\right$. (3.16)
Theorem 10.
Suppose that (H-M) holds. Then the RGBDSDE-NL (3.16) has a unique solution $(Y,Z,K)\in\mathcal{B}^{2}_{\theta,\mu}$ for any $\theta,\mu>0$.
Before proving Theorem 10, we record the following remark, which is identical to [12, Remark 2.3]:
Remark 11.
Let $(Y,Z,K)$ be a solution of the RGBDSDE-NL (1.5) associated with $(\xi,f,h,\kappa,g,S)$. For any $\mu>0$, we define
$\begin{split}\hat{Y}_{t}:=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}Y_{t},\quad\hat{Z}_{t}:=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}Z_{t},\quad d\hat{K}_{t}:=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}dK_{t}.\end{split}$
Using Corollary 7 with $\Psi(x,y,z)=e^{x+\mu y}z$, $\mathsf{A}_{t}=\left(\int_{0}^{t}\lambda_{s}\,ds,\kappa_{t}\right)$, and $\mathcal{Y}=Y$, we have
[…]
$\hat{Y}_{t}\geq\hat{S}_{t}$ for any $t\in[0,T]$ and $\int_{0}^{T}\left(\hat{Y}_{s}-\hat{S}_{s}\right)d\hat{K}_{s}=0$, with
$\left\{\begin{split}\hat{\xi}&=e^{\int_{0}^{T}\lambda_{s}ds+\mu\kappa_{T}}\xi,\\ \hat{f}(t,y,z)&=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}f\left(t,e^{-\int_{0}^{t}\lambda_{u}du-\mu\kappa_{t}}y,e^{-\int_{0}^{t}\lambda_{u}du-\mu\kappa_{t}}z\right)-\lambda_{t}y,\\ \hat{h}(t,y)&=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}h\left(s,e^{-\int_{0}^{t}\lambda_{u}du-\mu\kappa_{t}}y\right)-\mu y,\\ \hat{g}(t,y,z)&=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}g\left(t,e^{-\int_{0}^{t}\lambda_{u}du-\mu\kappa_{t}}y,e^{-\int_{0}^{s}\lambda_{u}du-\mu\kappa_{t}}z\right),\\ \hat{S}_{t}&=e^{\int_{0}^{t}\lambda_{s}ds+\mu\kappa_{t}}S_{t}.\end{split}\right$.
[…]
Proof of Theorem 10.
Uniqueness.
The uniqueness of the solution is established in Proposition 9.
Existence.
For the existence, the proof is divided into two main parts.
Part I:
In this part, we suppose that there exists a constant $M>0$ such that
$\sup_{0\leq t\leq T}\left|\varphi_{t}\right|^{2}+\sup_{0\leq t\leq T}\left|\psi_{t}\right|^{2}\leq M.$ (3.17)
Since $f$ is independent of $z$ and $g$ is independent of $(y,z)$, we set $\mathfrak{f}(t,y):=f(t,y,z)$ and $\mathfrak{g}(t):=g(t,y,z)$ for all $(t,y,z)\in[0,T]\times\mathbb{R}\times\ell^{2}$. In view of Remark 11, assumptions (H-M) (ii) and (iii) can be rewritten as follows:
(H-M)
  • (ii’)
    For all $y$, $y^{\prime}\in\mathbb{R}$, $d\mathbb{P}\otimes dt$-a.e.,
    $\left(y-y^{\prime}\right)\left(\mathfrak{f}\left(t,y\right)-\mathfrak{f}\left(t,y^{\prime}\right)\right)\leq 0.$
  • (iii’)
    For all $y$, $y^{\prime}\in\mathbb{R}$, $d\mathbb{P}\otimes d\kappa_{t}$-a.e.,
    $\left(y-y^{\prime}\right)\left(h\left(t,y\right)-h\left(t,y^{\prime}\right)\right)\leq 0.$
Step 1: Yosida approximation of the RGBDSDE-NL (1.5).
From (H-M)-(ii’)-(iii’) and [37, Annex B, p. 524], it follows that for every, $\left(\omega,t,y\right)\in\Omega\times[0,T]\times\mathbb{R}$ and $\delta>0$, there exists a unique $J_{\delta}^{\mathfrak{f}}=J_{\delta}^{\mathfrak{f}}(\omega,t,y)$, $J_{\delta}^{h}=J_{\delta}^{h}(\omega,t,y)\in\mathbb{R}$ such that
$J_{\delta}^{\mathfrak{f}}-\delta\mathfrak{f}(\omega,t,J^{\mathfrak{f}}_{\delta})=y,\quad J_{\delta}^{h}-\delta h(\omega,t,J^{h}_{\delta})=y.$
The Yosida approximation of $\mathfrak{f}$ and $h$ is defined respectively by $\mathfrak{f}_{\delta}=\mathfrak{f}_{\delta}\left(\omega,t,y\right)$ and $h_{\delta}=h_{\delta}\left(\omega,t,y\right)\in\mathbb{R}$ such that
$\begin{split}&\mathfrak{f}_{\delta}(\omega,t,y):=\dfrac{1}{\delta}\left(J_{\delta}^{\mathfrak{f}}(\omega,t,y)-y\right)=\mathfrak{f}\left(t,y+\delta\mathfrak{f}_{\delta}(\omega,t,y)\right),\\ &h_{\delta}(\omega,t,y):=\dfrac{1}{\delta}\left(J_{\delta}^{h}(\omega,t,y)-y\right)=h\left(t,y+\delta h_{\delta}(\omega,t,y)\right).\end{split}$ (3.18)
Note that $(\mathfrak{f}_{\delta},h_{\delta})$ is the unique triplet satisfying the system (3.18).
From [37, Annex B, Proposition 6.7], recall that for every $y\in\mathbb{R}$, the processes $\mathfrak{f}_{\delta}(\cdot,\cdot,y)$ and $h_{\delta}(\cdot,\cdot,y):\Omega\times[0,T]\rightarrow\mathbb{R}$ are jointly measurable and $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$. Furthermore, it holds that:
  • (Y)
    $\forall\delta,\mathsf{r}>0$, $\forall t\in[0,T]$, $\forall y,y^{\prime}\in\mathbb{R}$, $\mathbb{P}$-a.s.
    • (i)
      $\left(y-y^{\prime}\right)\left(\mathfrak{f}_{\delta}(t,y)-\mathfrak{f}_{\delta}(t,y^{\prime})\right)\leq 0$.
    • (ii)
      $\left(y-y^{\prime}\right)\left(h_{\delta}(t,y)-h_{\delta}(t,y^{\prime})\right)\leq 0$.
    • (iii)
      $\left|\mathfrak{f}_{\delta}(t,y)-\mathfrak{f}_{\delta}\left(t,y^{\prime}\right)\right|+\left|h_{\delta}(t,y)-h_{\delta}(t,y^{\prime})\right|\leq\dfrac{2}{\delta}\left|y-y^{\prime}\right|$.
    • (iv)
      $\left|\mathfrak{f}_{\delta}(t,y)\right|\leq\left|\mathfrak{f}(t,y)\right|$ and $\lim\limits_{\delta\rightarrow 0}\mathfrak{f}_{\delta}(t,y)=\mathfrak{f}(t,y)$,
      $\left|h_{\delta}(t,y)\right|\leq\left|h(t,y)\right|$ and $\lim\limits_{\delta\rightarrow 0}h_{\delta}(t,y)=h(t,y)$.
    • (v)
      $\left(y-y^{\prime}\right)\left(\mathfrak{f}_{\delta}(t,y)-\mathfrak{f}_{\mathsf{r}}(t,y^{\prime})\right)\leq\left(\delta+\mathsf{r}\right)\mathfrak{f}_{\delta}(t,y)\mathfrak{f}_{\mathsf{r}}(t,y^{\prime})$,
      $\left(y-y^{\prime}\right)\left(h_{\delta}(t,y)-h_{\mathsf{r}}(t,y^{\prime})\right)\leq\left(\delta+\mathsf{r}\right)h_{\delta}(t,y)h_{\mathsf{r}}(t,y^{\prime})$.
Let $0<\delta\leq 1$. From Theorem 34, there exists a unique solution $\left(Y^{\delta},Z^{\delta},K^{\delta}\right)\in\left(\mathcal{S}^{2}_{\mu}\cap\mathcal{H}^{2,\mathcal{Q}}_{\mu}\right)\times\mathcal{H}^{2,\ell^{2}}_{\mu}\times\mathcal{A}^{2}$ of the RGBDSDE-NL (3.16) associated with $\left({\xi},\mathfrak{f}_{\delta},h_{\delta},\kappa,\mathfrak{g},S\right)$. In other words, the approximating equation:
$\left\{\begin{split}\text{(i)}&~Y^{\delta}_{t}={\xi}+\int_{t}^{T}{\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})ds+\int_{t}^{T}{h}_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}+\int_{t}^{T}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\left(K^{\delta}_{T}-K^{\delta}_{t}\right)-\sum_{k=1}^{d}\int_{t}^{T}Z^{\delta,(k)}_{s}dH^{(k)}_{s},\\ \text{(ii)}&~Y^{\delta}_{t}\geq S_{t},~t\in[0,T]~\mbox{ and }~\int_{0}^{T}\left(Y^{\delta}_{s}-S_{s}\right)dK^{\delta}_{s}=0,\end{split}\right$. (3.19)
has a unique solution $\left(Y^{\delta},Z^{\delta},K^{\delta}\right)\in\left(\mathcal{S}^{2}_{\mu}\cap\mathcal{H}^{2,\mathcal{Q}}_{\mu}\right)\times\mathcal{H}^{2,\ell^{2}}_{\mu}\times\mathcal{A}^{2}$.
Step 2: Uniform estimation of $\left\{\left(Y^{\delta},Z^{\delta},K^{\delta}\right)\right\}_{0<\delta\leq 1}$.
We aim to show the following lemma:
Lemma 12.
For any $0<\delta\leq 1$ and for every $\theta,\mu>0$, the process $\left(Y^{\delta},Z^{\delta},K^{\delta}\right)$ satisfies
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\mathbb{E}\left|K^{\delta}_{T}\right|^{2}\\ &\leq\mathfrak{c}_{\theta,\mu,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}\left(S_{t}\right)^{+}\right|^{2}\right)\end{split}$
Proof.
Using Corollary 7, we derive that
$\begin{split}&\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}+\theta\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mu\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}d\kappa_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &=\Phi^{\theta,\mu}_{T}\left|{\xi}\right|^{2}+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y^{\delta}_{s}{\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})ds+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y^{\delta}_{s}{h}_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}\\ &\quad+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y^{\delta}_{s}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds-2\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y^{\delta}_{s-}Z^{\delta,(k)}_{s}dH^{(k)}_{s}\\ &\quad-\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}Z_{s}^{\delta,(k)}Z_{s}^{\delta,(k^{\prime})}\,\left(d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}-d\big\langle H^{(k)},H^{(k^{\prime})}\big\rangle_{s}\right)+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y^{\delta}_{s}dK^{\delta}_{s}.\end{split}$ (3.20)
Using (Y)-(i)-(ii)-(iv) and the basic inequality $2ab\leq\varepsilon a^{2}+\frac{1}{\varepsilon}b^{2}$ for any $\varepsilon>0$, we derive
$\begin{split}2Y^{\delta}_{s}{\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})ds\leq 2Y^{\delta}_{s}{\mathfrak{f}}_{\delta}(s,0)ds&\leq 2\left|Y^{\delta}_{s}\right|\left|{\mathfrak{f}}(s,0)\right|ds\leq\varepsilon^{\prime}\left|Y^{\delta}_{s}\right|^{2}dV_{s}+\frac{1}{\varepsilon^{\prime}}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds,\end{split}$
and
$\begin{split}2Y^{\delta}_{s}h_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}\leq 2Y^{\delta}_{s}h_{\delta}(s,0)d\kappa_{s}&\leq 2\left|Y^{\delta}_{s}\right|\left|h(s,0)\right|d\kappa_{s}\leq\varepsilon"\left|Y^{\delta}_{s}\right|^{2}d\kappa_{s}+\frac{1}{\varepsilon"}\left|\psi_{s}\right|^{2}d\kappa_{s},\end{split}$
for any $\varepsilon^{\prime},\varepsilon>0$.
Then, plugging these inequalities into (3.20), along with the Skorokhod condition and taking the expectation, we get
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}+\left(\theta-\varepsilon^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+(\mu-\varepsilon")\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathbb{E}\Phi^{\theta,\mu}_{T}\left|{\xi}\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}S_{s}dK^{\delta}_{s}.\end{split}$ (3.21)
Choosing $\varepsilon^{\prime}<\theta$ and $\varepsilon^{\prime\prime}<\mu$, we derive, for any $\varepsilon^{\prime\prime\prime}>0$
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds\right)\\ &\qquad+\varepsilon^{\prime\prime\prime}\mathfrak{c}_{\theta,\mu}^{2}\mathbb{E}\sup_{t\leq u\leq T}\left|\Phi^{\theta,\mu}_{u}(S_{u})^{+}\right|^{2}+\frac{1}{\varepsilon^{\prime\prime\prime}}\mathbb{E}\left|K^{\delta}_{T}-K^{\delta}_{t}\right|^{2}.\end{split}$ (3.22)
Returning to (1.5)-(i), by squaring, we have
$\begin{split}&\mathbb{E}\left|K^{\delta}_{T}-K^{\delta}_{t}\right|^{2}\\ &\leq 6\left(\mathbb{E}\left|Y^{\delta}_{t}\right|^{2}+\mathbb{E}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}\right)^{2}\right.\\ &\left.+\mathbb{E}\left(\int_{t}^{t}\mathfrak{g}(s)\overleftarrow{dB}_{s}\right)^{2}+\mathbb{E}\left(\sum_{k=1}^{d}\int_{t}^{T}Z^{\delta,(k)}_{s}dH^{(k)}_{s}\right)^{2}\right)\\ &\leq 6\left(\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}+\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}\right)^{2}\right.\\ &\qquad\left.+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right)\end{split}$ (3.23)
Using (Y)-(iv), (H-M)-(ix), and Hölder’s inequality, we have
$\begin{split}\mathbb{E}\left(\int_{t}^{T}\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})ds\right)^{2}&\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{-\theta,\mu}_{s}dV_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})}{a_{s}}\right|^{2}ds\right)\\ &\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{-\theta,\mu}_{s}dV_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\mathfrak{f}(s,Y^{\delta}_{s})}{a_{s}}\right|^{2}ds\right)\\ &\leq\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\frac{\varphi^{2}_{s}+\phi_{s}^{2}\left|Y^{\delta}_{s}\right|^{2}}{a^{2}_{s}}ds\\ &\leq\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}dV_{s}+\frac{3}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds,\end{split}$ (3.24)
and
$\begin{split}\mathbb{E}\left(\int_{t}^{T}h_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}\right)^{2}&\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{\theta,-\mu}_{s}d\kappa_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|h_{\delta}(s,Y^{\delta}_{s})\right|^{2}d\kappa_{s}\right)\\ &\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{\theta,-\mu}_{s}d\kappa_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|h(s,Y^{\delta}_{s})\right|^{2}d\kappa_{s}\right)\\ &\leq\frac{2\zeta^{2}}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}d\kappa_{s}+\frac{2}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\psi^{2}_{s}d\kappa_{s}.\end{split}$ (3.25)
Substituting this into (3.22), and choosing $\varepsilon^{\prime\prime\prime}>\frac{18}{\theta}+\frac{12\zeta^{2}}{\mu}+12$, we have
$\begin{split}&\sup_{0\leq t\leq T}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta}_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$ (3.26)
Then, from (3.23), we conclude
$\begin{split}\mathbb{E}\left|K^{\delta}_{T}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$
The remainder of the proof relies on the uniform estimation (3.26) and the Burkholder-Davis-Gundy inequality. Using a similar computation as that in Proposition 8, we can easily derive that, for any $\theta,\mu>0$,
$\begin{split}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$
Completing the proof of Lemma 12. ∎
Step 3: The process $\left\{Y^{\delta},Z^{\delta}\right\}_{0<\delta\leq 1}$ is a Cauchy sequence in $\mathcal{C}^{2}_{\theta,\mu}\times\mathcal{H}^{2,\ell}_{\theta,\mu}$.
Let $0<\delta,\delta^{\prime}\leq 1$. Set $\mathcal{R}^{\delta,\delta^{\prime}}=\mathcal{R}^{\delta}-\mathcal{R}^{\delta^{\prime}}$ with $\mathcal{R}\in\{Y,Z,K\}$. Using Corollary 7 with expectation, we arrive at
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y^{\delta,\delta^{\prime}}_{t}\right|^{2}+\theta\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta,\delta^{\prime}}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\delta,\delta^{\prime}}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z^{\delta}_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &=2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)\left({\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})-{\mathfrak{f}}_{\delta^{\prime}}(s,Y^{\delta^{\prime}}_{s})\right)ds\\ &\quad+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)\left({h}_{\delta}(s,Y^{\delta}_{s})-{h}_{\delta^{\prime}}(s,Y^{\delta^{\prime}}_{s})\right)d\kappa_{s}+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)dK^{\delta,\delta^{\prime}}_{s}.\end{split}$ (3.27)
Using the Skorokhod condition (3.19)-(ii), we obtain
$\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)dK^{\delta,\delta^{\prime}}_{s}=-\left(Y^{\delta}_{s}-S_{s}\right)dK^{\delta^{\prime}}_{s}+\left(S_{s}-Y^{\delta^{\prime}}_{s}\right)dK^{\delta}_{s}\leq 0.$
Next, from (Y)-(iv)-(v), we can deduce
$\begin{split}&2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)\left({\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})-{\mathfrak{f}}_{\delta^{\prime}}(s,Y^{\delta^{\prime}}_{s})\right)ds\\ &\leq 2\left(\delta+\delta^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})\right|\left|\mathfrak{f}_{\delta^{\prime}}(s,Y^{\delta^{\prime}}_{s})\right|ds\\ &\leq 2\left(\delta+\delta^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{f}(s,Y^{\delta}_{s})\right|\left|\mathfrak{f}(s,Y^{\delta^{\prime}}_{s})\right|ds\\ &\leq 2\left(\delta+\delta^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(\varphi_{s}+\phi_{s}\left|Y^{\delta}_{s}\right|\right)\left(\varphi_{s}+\phi_{s}\left|Y^{\delta^{\prime}}_{s}\right|\right)ds\\ &=2\left(\delta+\delta^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(\varphi^{2}_{s}+\varphi_{s}\phi_{s}\left|Y^{\delta}_{s}\right|+\varphi_{s}\phi_{s}\left|Y^{\delta^{\prime}}_{s}\right|+\phi_{s}^{2}\left|Y^{\delta^{\prime}}_{s}\right|\left|Y^{\delta^{\prime}}_{s}\right|\right)ds\\ &\leq\left(\delta+\delta^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\{2M^{2}ds+\left(2M+1\right)\left\{\left|Y^{\delta^{\prime}}_{s}\right|^{2}+\left|Y^{\delta^{\prime}}_{s}\right|^{2}\right\}dV_{s}\right\}\\ &\leq\frac{2M^{2}\left(\delta+\delta^{\prime}\right)}{\epsilon}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}d\mathcal{Q}_{s}+\left(2M+1\right)\left(\delta+\delta^{\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\{\left|Y^{\delta^{\prime}}_{s}\right|^{2}+\left|Y^{\delta^{\prime}}_{s}\right|^{2}\right\}dV_{s}\\ &\leq\left(\delta+\delta^{\prime}\right)\mathfrak{c}_{M,\epsilon}\mathbb{E}\int_{t}^{T}\left(\Phi^{\theta,\mu}_{s}d\mathcal{Q}_{s}+\left\{\left|Y^{\delta^{\prime}}_{s}\right|^{2}+\left|Y^{\delta^{\prime}}_{s}\right|^{2}\right\}dV_{s}\right).\end{split}$
Using Lemma 12 and Remark 5, we conclude that there exists a constant $\mathfrak{c}_{M,\epsilon}$ (independent of $\delta$ and $\delta^{\prime}$) such that for any $t\in[0,T]$, we have
$\begin{split}&2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)\left({\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})-{\mathfrak{f}}_{\delta^{\prime}}(s,Y^{\delta^{\prime}}_{s})\right)ds\leq\left(\delta+\delta^{\prime}\right)\mathfrak{c}_{M}.\end{split}$
Similarly, we can show that
$\begin{split}&2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(Y^{\delta}_{s}-Y^{\delta^{\prime}}_{s}\right)\left(h_{\delta}(s,Y^{\delta}_{s})-h_{\delta^{\prime}}(s,Y^{\delta^{\prime}}_{s})\right)d\kappa_{s}\leq\left(\delta+\delta^{\prime}\right)\mathfrak{c}_{M,\zeta}.\end{split}$
By the dominated convergence theorem, it follows that the right-hand side of (3.27) goes 0 when $\delta,\delta^{\prime}\rightarrow 0^{+}$. This implies that, for any $\theta,\mu>0$
$\lim\limits_{\delta,\delta^{\prime}\rightarrow 0^{+}}\left(\big\|Y^{\delta}-Y^{\delta^{\prime}}\big\|^{2}_{\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}}+\big\|Z^{\delta}-Z^{\delta^{\prime}}\big\|^{2}_{\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}}\right)=0.$ (3.28)
Now, using the Burkholder-Davis-Gundy inequality and following a similar computation as in Proposition 8 (or 9), along with the convergence result (3.28), we can derive that
$\lim\limits_{\delta,\delta^{\prime}\rightarrow 0^{+}}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y^{\delta}_{t}-Y^{\delta^{\prime}}_{t}\right|^{2}=0.$
Then, $\left\{Y^{\delta},Z^{\delta}\right\}_{0<\delta\leq 1}$ converges in $\mathcal{C}^{2}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$, i.e., there exists a process $(Y,Z)$ such that $Y^{\delta}\rightarrow Y$ in $\mathcal{C}^{2}_{\theta,\mu}$ and $Z^{\delta}\rightarrow Z$ in $\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$ as $\delta\rightarrow 0^{+}$.
Step 4: Convergence of the process $\{K^{\delta}\}_{0<\delta\leq 1}$ in $\mathcal{A}^{2}$ and the integrals $\{\int_{0}^{\cdot}\mathfrak{f}(s,Y^{\delta}_{s})ds,\int_{0}^{\cdot}h(s,Y^{\delta}_{s})d\kappa_{s}\}_{0<\delta\leq 1}$ in $\mathbb{L}^{2}$.
From the RGBDSDE-NL (3.19), for any $0\leq\delta\leq 1$, we have
$K^{\delta}_{t}=Y^{\delta}_{0}-Y^{\delta}_{t}-\int_{0}^{t}{\mathfrak{f}}_{\delta}(s,Y^{\delta}_{s})ds-\int_{0}^{t}{h}_{\delta}(s,Y^{\delta}_{s})d\kappa_{s}-\int_{0}^{t}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\sum_{k=1}^{d}\int_{0}^{t}Z^{\delta,(k)}_{s}dH^{(k)}_{s}$.
By employing a similar argument, we also obtain
$\lim\limits_{\delta,\delta^{\prime}\rightarrow 0^{+}}\mathbb{E}\sup_{0\leq t\leq T}\left|K^{\delta}_{t}-K^{\delta^{\prime}}_{t}\right|^{2}=0.$ (3.29)
In conclusion $\left\{Y^{\delta},Z^{\delta},K^{\delta}\right\}$ is a Cauchy sequence in $\mathcal{B}^{2}_{\theta,\mu}$.
Now, from the definition of Yosida approximation (3.18), we have
$\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)=\mathfrak{f}\left(s,Y^{\delta}_{s}+\delta\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)\right)~\mbox{ and }~h_{\delta}\left(s,Y^{\delta}_{s}\right)=h\left(s,Y^{\delta}_{s}+\delta h_{\delta}\left(s,Y^{\delta}_{s}\right)\right).$
Using Lemma 12, along with conditions (Y)-(iv) and (H-M)-(ix), we have
$\begin{split}\mathbb{E}\left(\int_{0}^{T}\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)ds\right)^{2}&\leq\frac{2}{\theta}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}\left\{\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\left|Y^{\delta}_{s}\right|^{2}dV_{s}\right\}\leq\mathfrak{c}_{\theta},\end{split}$ (3.30)
and
$\begin{split}\mathbb{E}\left(\int_{0}^{T}h_{\delta}\left(s,Y^{\delta}_{s}\right)d\kappa_{s}\right)^{2}&\leq\frac{2}{\mu}\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}\left\{\psi^{2}_{s}ds+\left|Y^{\delta}_{s}\right|^{2}d\kappa_{s}\right\}\leq\mathfrak{c}_{\mu,\zeta}.\end{split}$ (3.31)
Then, for any $t\in[0,T]$, we have $\lim_{\delta\rightarrow 0^{+}}\int_{0}^{t}\delta\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})\,ds=0$ in $\mathbb{L}^{2}$ and $\lim_{\delta\rightarrow 0^{+}}\int_{0}^{t}\delta h_{\delta}(s,Y^{\delta}_{s})\,d\kappa_{s}=0$ in $\mathbb{L}^{2}$. Therefore, by applying the partial converse of the Dominated Convergence Theorem, we deduce the existence of two subsequences $\left(\delta_{k}\mathfrak{f}_{\delta_{k}}(s,Y^{\delta_{k}}_{s})\right)_{k\in\mathbb{N}}$ and $\left(\delta_{k}h_{\delta_{k}}(s,Y^{\delta_{k}}_{s})\right)_{k\in\mathbb{N}}$ such that $\lim_{k\rightarrow+\infty}\delta_{k}\mathfrak{f}_{\delta_{k}}(s,Y^{\delta_{k}}_{s})=0$ $d\mathbb{P}\otimes ds$-a.e. and $\lim_{k\rightarrow+\infty}\delta_{k}h_{\delta_{k}}(s,Y^{\delta_{k}}_{s})=0$ $d\mathbb{P}\otimes d\kappa_{s}$-a.e. For simplicity, we continue to denote these two subsequences by the original notation, thus we have $\lim_{\delta\rightarrow 0^{+}}\delta\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})=0$ $d\mathbb{P}\otimes ds$-a.e. and $\lim_{\delta\rightarrow 0^{+}}\delta h_{\delta}(s,Y^{\delta}_{s})=0$ $d\mathbb{P}\otimes d\kappa_{s}$-a.e. Moreover, given that $\left\{Y^{\delta}\right\}_{0<\delta\leq 1}$ converges in $\mathcal{S}^{2}_{\theta,\mu}$ to $Y$, we have $\lim_{\delta\rightarrow 0^{+}}Y^{\delta}_{s}=Y_{s}$ a.s. for any $t\in[0,T]$. This, together with the continuity of the functions $y\mapsto\mathfrak{f}(s,y)$ and $y\mapsto h(s,y)$, allows us to deduce that
$\lim\limits_{\delta\rightarrow 0^{+}}\mathfrak{f}_{\delta}\left(s,Y^{\delta}_{s}\right)=\mathfrak{f}\left(s,Y_{s}\right)~d\mathbb{P}\otimes ds\mbox{-a.e.}\mbox{ and }\lim\limits_{\delta\rightarrow 0^{+}}h_{\delta}\left(s,Y^{\delta}_{s}\right)=h\left(s,Y_{s}\right)~d\mathbb{P}\otimes d\kappa_{s}\mbox{-a.e.}$
Combining this with (3.30) and (3.31), and applying the Lebesgue Dominated Convergence Theorem, we obtain $\lim_{\delta\rightarrow 0^{+}}\int_{\cdot}^{T}\mathfrak{f}_{\delta}(s,Y^{\delta}_{s})\,ds=\int_{\cdot}^{T}\mathfrak{f}(s,Y_{s})\,ds$ and $\lim_{\delta\rightarrow 0^{+}}\int_{\cdot}^{T}h_{\delta}(s,Y^{\delta}_{s})\,d\kappa_{s}=\int_{t}^{T}h(s,Y_{s})$ in $\mathcal{S}^{2}$.
Step 5: The limiting process $(Y,Z,K)$ satisfies the RGBDSDE-NL (3.16).
Using the Burkholder-Davis-Gundy inequality and (3.28), we derive that $\lim_{\delta\rightarrow 0^{+}}\sum_{k=1}^{d}\int_{\cdot}^{T}Z^{\delta,(k)}_{s}\,dH^{(k)}_{s}=\sum_{k=1}^{d}\int_{\cdot}^{T}Z^{(k)}_{s}\,dH^{(k)}_{s}$ in $\mathcal{S}^{2}_{\theta,\mu}$. Then, taking the limit term by term in $\mathbb{L}^{2}$ in (3.19) along with the obtained convergence result, we get
$Y_{t}=\xi+\int_{t}^{T}\mathfrak{f}(s,Y_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}+\int_{t}^{T}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\left(K_{T}-K_{t}\right)-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s}$.
Additionally, from (3.19-(ii)), we have $Y_{t}\geq S_{t}$ for any $t\in[0,T]$. Finally, from (3.29), it follows that $\lim_{\delta\rightarrow 0^{+}}\|K^{\delta}-K\|^{2}_{\mathcal{A}^{2}}=0$. Thus, the sequence $\{K^{\delta}\}_{0<\delta\leq 1}$ converges to $K$ uniformly in $t$ in probability. Consequently, the measure $dK^{\delta}$ converges to $dK$ weakly in probability, leading to $\lim_{\delta\rightarrow 0^{+}}\int_{0}^{T}(Y^{\delta}_{s}-S_{s})\,dK^{\delta}_{s}=\int_{0}^{T}(Y_{s}-S_{s})\,dK_{s}=0$ in probability. Passing again to a subsequence when necessary, we derive that $\int_{0}^{T}(Y_{s}-S_{s})\,dK_{s}=0$ almost surely. Note that Lemma 12, Fatou’s Lemma, and the convergence results allow us to derive that the limiting process verifies
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\mathbb{E}\left|K_{T}\right|^{2}\\ &\leq\mathfrak{c}_{\theta,\mu,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\mathfrak{g}(s)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}\left(S_{t}\right)^{+}\right|^{2}\right).\end{split}$
Part II: Beyond the boundedness condition (3.17)
For any $\mathsf{p}\geq 1$, we consider
[…]
Condition (3.17) is satisfied:
$\sup_{0\leq t\leq T}\left|\mathfrak{f}_{\mathsf{p}}(t,0)\right|^{2}+\sup_{0\leq t\leq T}\left|h_{\mathsf{p}}(t,0)\right|^{2}\leq\mathfrak{c}_{p}.$
By Part I, there exists a unique process $\left(Y^{\mathsf{p}},Z^{\mathsf{p}},K^{\mathsf{p}}\right)\in\mathcal{B}^{2}_{\theta,\mu}$ such that
$\left\{\begin{split}\text{(i)}&~Y^{\mathsf{p}}_{t}={\xi}+\int_{t}^{T}{\mathfrak{f}}_{\mathsf{p}}(s,Y_{s})ds+\int_{t}^{T}{h}_{\mathsf{p}}(s,Y_{s})d\kappa_{s}+\int_{t}^{T}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\left(K^{p}_{T}-K^{p}_{t}\right)-\sum_{k=1}^{d}\int_{t}^{T}Z^{\mathsf{p},(k)}_{s}dH^{(k)}_{s},\\ \text{(ii)}&~Y^{\mathsf{p}}_{t}\geq S_{t},~t\in[0,T]~\mbox{ and }~\int_{0}^{T}\left(Y^{\mathsf{p}}_{s}-S_{s}\right)dK^{\mathsf{p}}_{s}=0.\end{split}\right$. (3.33)
From Definition (3.32) and (H-M)-(xiii), we have
$\begin{split}\lim\limits_{\mathsf{p}\rightarrow+\infty}\left(\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\mathfrak{f}_{\mathsf{p}}(t,0)-\mathfrak{f}(t,0)}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|h_{\mathsf{p}}(t,0)-h(t,0)\right|^{2}d\kappa_{s}\right)=0.\end{split}$ (3.34)
Now let $\mathsf{q}>\mathsf{p}$. By using Itô’s formula, the Skorokhod condition in (3.33), the Burkholder-Davis-Gundy inequality, and familiar computations, we end up with
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y^{\mathsf{q}}_{t}-Y^{\mathsf{p}}_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y^{\mathsf{q}}_{s}-Y^{\mathsf{p}}_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|\left(Z^{\mathsf{q}}_{s}-Z^{\mathsf{p}}_{s}\right)\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\zeta}\left(\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\mathfrak{f}_{\mathsf{q}}(s,0)-\mathfrak{f}_{\mathsf{p}}(s,0)}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|h_{\mathsf{q}}(s,0)-h_{\mathsf{p}}(s,0)\right|^{2}d\kappa_{s}\right).\end{split}$ (3.35)
This, along with (3.34), allows us to conclude that the right-hand side of (3.35) tends to 0 as $\mathsf{p},\mathsf{q}\rightarrow+\infty$. In addition, for any $\mathsf{p}\geq 1$,
$K^{\mathsf{p}}_{t}=Y^{\mathsf{p}}_{0}-Y^{\mathsf{p}}_{t}-\int_{0}^{t}{\mathfrak{f}}_{\mathsf{p}}(s,Y^{\mathsf{p}}_{s})ds-\int_{0}^{t}{h}_{\mathsf{p}}(s,Y^{\mathsf{p}}_{s})d\kappa_{s}-\int_{0}^{t}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\sum_{k=1}^{d}\int_{0}^{t}Z^{\mathsf{p},(k)}_{s}dH^{(k)}_{s}$.
Then, we also get
$\lim\limits_{\mathsf{p},\mathsf{q}\rightarrow+\infty}\mathbb{E}\sup_{0\leq t\leq T}\left|K^{\mathsf{p}}_{t}-K^{\mathsf{q}}_{t}\right|^{2}=0.$
Similarly to Part 1, we derive that $\left\{\left(Y^{\mathsf{p}},Z^{\mathsf{p}},K^{\mathsf{p}}\right)\right\}_{\mathsf{p}\geq 1}$ is a Cauchy sequence in $\mathcal{B}^{2}_{\beta,\mu}$, and the corresponding limit process $(Y,Z,K)\in\mathcal{B}^{2}_{\beta,\mu}$ verifies the RGBDSDE-NL (3.16). ∎
3.3.2 General case
Here, the drivers are taken in full generality, with $f$ depending on $z$ and $g$ on $(y,z)$. We now present the section’s main result.
Theorem 13.
Suppose that (H-M) holds. Then, there exists a constant $\theta_{0}>0$ such that the RGBDSDE-NL (1.5) has a unique solution $\left(Y,Z,K\right)\in\mathcal{B}^{2}_{\theta,\mu}$ for any $\theta\geq\theta_{0}$ and $\mu>0$.
Uniqueness.
The uniqueness has previously been proven by Proposition 9.
Existence.
Using the Picard approximation sequence as our method for proving existence, we examine the sequence $\left\{Y^{n},Z^{n},K^{n}\right\}_{n\geq 0}\subset\mathcal{B}^{2}_{\theta,\mu}$ defined recursively as follows: $Y^{0}=Z^{0}=K^{0}=0$. Then, for any $n\geq 0$, we let $(Y^{n+1},Z^{n+1},K^{n+1})$ be the unique solution of the RGBDSDE-NL (1.5) associated with $(\xi,f(t,Y^{n+1}_{t},Z^{n}_{t}),h(t,Y^{n+1}_{t}),\kappa_{t},g(t,Y^{n}_{t},Z^{n}_{t}),S_{t})$. In other words, the process $(Y^{n+1},Z^{n+1},K^{n+1})$ satisfies
[…]
Using Theorem 10, we deduce that for every $n\geq 1$, the RGBDSDE-NL (3.36) has a unique solution. To simplify notations, we set $\hat{\mathcal{R}}^{n+1}=\mathcal{R}^{n+1}-\mathcal{R}^{n}$ for $\mathcal{R}^{n}\in\left\{Y^{n},Z^{n},K^{n}\right\}$ and define $\mathfrak{f}_{n}(t)=f(t,Y^{n}_{t},Z^{n-1}_{t})-f(t,Y^{n-1}_{t},Z^{n-2}_{t})$, $\mathfrak{h}_{n}(t)=h(t,Y^{n}_{t})-h(t,Y^{n-1}_{t})$, and $\mathfrak{g}_{n}(t)=g(t,Y^{n}_{t},Z^{n}_{t})-g(t,Y^{n-1}_{t},Z^{n-1}_{t})$. From (3.36)-(ii), the state process $\hat{Y}^{n+1}$ satisfies the following BDSDE:
$\begin{split}\hat{Y}^{n+1}_{t}=&\int_{t}^{T}\mathfrak{f}_{n+1}(s)ds+\int_{t}^{T}\mathfrak{h}_{n+1}(s)dA_{s}+\int_{t}^{T}\mathfrak{g}_{n}(s)\overleftarrow{dB}_{s}\\ &+\left(\hat{K}^{n+1}_{T}-\hat{K}^{n+1}_{t}\right)-\sum_{k=1}^{d}\int_{t}^{T}Z^{n+1,(k)}_{s}dH^{(k)}_{s},\quad t\in[0,T].\end{split}$ (3.37)
To apply Corollary 7 to the dynamic equation (3.37), we require the following estimations concerning the newly introduced generators. Utilizing conditions (H-M)-(ii), (iii), (v), and (vii), we obtain the following results for any $\varepsilon_{1},\varepsilon_{2}>0$:
$\begin{split}2\hat{Y}^{n+1}_{s}\mathfrak{f}_{n+1}(s)ds&\leq 2\left|\hat{Y}^{n+1}_{s}\right|\left(\lambda_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}+\eta_{s}\left\|\hat{Z}^{n}_{s}\gamma_{s}\right\|_{\ell^{2}}\right)ds\\ &\leq\left(2+\frac{1}{\varepsilon_{1}}\right)\left|\hat{Y}^{n+1}_{s}\right|^{2}dV_{s}+\varepsilon_{1}\left\|\gamma_{s}\hat{Z}^{n}_{s}\right\|^{2}_{\ell^{2}}ds,\end{split}$ (3.38)
and
$\begin{split}2\hat{Y}^{n+1}_{s}\mathfrak{h}_{n+1}{\hat{\Theta}^{n}}(s)dA_{s}&\leq 2\varrho_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}d\kappa_{s}\leq\varepsilon_{2}\left|\hat{Y}^{n}_{s}\right|^{2}d\kappa_{s}.\end{split}$ (3.39)
and
$\begin{split}\left|\mathfrak{g}_{n}(s)\right|^{2}ds&\leq\rho_{s}\left|\hat{Y}^{n}_{s}\right|^{2}ds+\alpha\left\|\gamma_{s}\hat{Z}^{n}_{s}\right\|^{2}_{\ell^{2}}ds\leq\left|\hat{Y}^{n}_{s}\right|^{2}dV_{s}+\alpha\left\|\gamma_{s}\hat{Z}^{n}_{s}\right\|^{2}_{\ell^{2}}ds.\end{split}$ (3.40)
Next, we apply the formula from Corollary 7 to ${\Phi}^{\theta,\mu}\left|\hat{Y}^{n+1}\right|^{2}$, where $\hat{Y}^{n+1}$ is defined by (3.37). Utilizing the Skorokhod condition of the reflection processes $\hat{K}^{n+1}$, which implies that $\hat{Y}^{n+1}_{s}d\hat{K}^{n+1}_{s}\leq 0$, along with inequalities (3.38), (3.39), (3.40), and following similar arguments as in Proposition 8, we derive, for any $\theta$, $\mu$, $\varepsilon_{1}$, $\varepsilon_{2}>0$,
$\begin{split}&\left(\theta-2-\frac{1}{\varepsilon_{1}}\right)\mathbb{E}\int_{t}^{T}{\Phi}^{\theta,\mu}_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}{\Phi}^{\theta,\mu}_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left\|\gamma_{s}\hat{Z}^{n+1}_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n}_{s}\right|^{2}dV_{s}+\varepsilon_{2}\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n}_{s}\right|^{2}d\kappa_{s}+\left(\varepsilon_{1}+\alpha\right)\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left\|\gamma_{s}\hat{Z}^{n}_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq(\varepsilon_{1}+\alpha)\left(\frac{2}{\varepsilon_{1}+\alpha}\mathbb{E}\int_{t}^{T}\hat{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n}_{s}\right|^{2}dV_{s}+\frac{\varepsilon_{2}}{\varepsilon_{1}+\alpha}\mathbb{E}\int_{t}^{T}\hat{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\hat{\Phi}^{\beta,\mu}_{s}\left\|\gamma_{s}\hat{Z}^{n}_{s}\right\|^{2}_{\ell^{2}}ds\right).\end{split}$
Fix $\varepsilon_{1}>0$, choose $\varepsilon_{2}=(\varepsilon_{1}+\alpha)\mu$, and define $\bar{\mathfrak{c}}=\frac{2}{\varepsilon_{1}+\alpha}$, $\theta_{0}=2+\frac{1}{\varepsilon_{1}}+\bar{\mathfrak{c}}$, and $\mu_{0}=1+\frac{1}{\varepsilon_{2}}$. By choosing $\theta\geq\theta_{0}$, we obtain
$\begin{split}&\bar{\mathfrak{c}}\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left\|\hat{Z}^{n+1}_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq(\varepsilon_{1}+\alpha)\left(\bar{\mathfrak{c}}\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left\|\hat{Z}^{n+1}_{s}\right\|^{2}_{\ell^{2}}ds\right).\end{split}$
Using simple iterations, we establish that
$\begin{split}&\bar{\mathfrak{c}}\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{n+1}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left\|\hat{Z}^{n+1}_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq(\varepsilon_{1}+\alpha)^{n}\left(\bar{\mathfrak{c}}\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{1}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left|\hat{Y}^{1}_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}{\Phi}^{\beta,\mu}_{s}\left\|\hat{Z}^{1}_{s}\right\|^{2}_{\ell^{2}}ds\right).\end{split}$ (3.41)
Choosing $\varepsilon_{1}>0$ such that $\varepsilon_{1}<1-\alpha$, we deduce that $\{Y^{n},Z^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$ for any $\theta\geq\theta_{0}$ and $\mu>0$. It remains to show that the sequence $\{Y^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{S}^{2}_{\theta,\mu}$. To this end, for any two integers $n,m\geq 0$, we set $\mathcal{R}^{n,m}=\mathcal{R}^{n}-\mathcal{R}^{m}$ for $\mathcal{R}^{n}\in\{Y^{n},Z^{n},K^{n}\}$. Then, we simply apply Corollary 7 to $\Phi^{\theta,\mu}_{\tau}|Y^{n+1,m+1}_{\tau}|^{2}$ along with the Burkholder-Davis-Gundy inequality, and we follow a similar argument as in Theorem 10, which yields
$\begin{split}&\mathbb{E}\left[\sup_{0\leq t\leq T}{\Phi}^{\beta,\mu}_{\tau}\left|Y^{n+1,m+1}_{t}\right|^{2}\right]\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha}\left(\left\|Y^{n+1,m+1}\right\|^{2}_{{\mathcal{H}}^{2,\mathcal{Q}}_{\theta,\mu}}+\left\|Z^{n+1,m+1}\right\|^{2}_{{\mathcal{H}}^{2,\ell^{2}}_{\theta,\mu}}+\left\|Y^{n,m}\right\|^{2}_{\mathcal{H}^{2,\mathcal{Q}}_{\beta,\mu}}+\left\|Z^{n,m}\right\|^{2}_{{\mathcal{H}}^{2,\ell^{2}}_{\theta,\mu}}\right).\end{split}$ (3.42)
Since $\{Y^{n},Z^{n},U^{n}\}_{n\geq 0}$ is a Cauchy sequence in the Banach space $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$ for any $\theta\geq\theta_{0}$ and $\mu>0$ (with $\theta_{0}$ defined above), we deduce that $\{Y^{n}\}_{n\geq 0}$ is also a Cauchy sequence in $\mathcal{S}^{2}_{\theta,\mu}$ for the same range of values of $\theta$ and $\mu$. Finally, by a familiar argument, we obtain
$\lim\limits_{n,m\rightarrow+\infty}\mathbb{E}\sup_{0\leq t\leq T}\left|{K}^{n+1,m+1}_{t}\right|^{2}=0.$
Therefore, the sequence $\{Y^{n},Z^{n},K^{n}\}_{n\geq 0}$ converges to a limit $(Y,Z,K)$ in the space $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}\times\mathcal{A}^{2}$ for any $\theta\geq\theta_{0}$ and $\mu>0$. In other words, the triplet $(Y,Z,K)$ satisfies
$\lim\limits_{n\rightarrow+\infty}\left\|\left(Y^{n}-Y,Z^{n}-Z,K^{n}-K\right)\right\|_{\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{{H}}^{2,\ell^{2}}_{\theta,\mu}\times\mathcal{A}^{2}}=0,\quad\forall(\beta,\mu)\in[\theta_{0},+\infty)\times(0,+\infty).$ (3.43)
Now, let’s focus on the convergence of the driver and martingale parts of (3.36)-(ii).
To this end, based on (3.43), we derive that $\{Y^{n},Z^{n}\}_{n\geq 0}$ is bounded in $\mathcal{H}^{2,\mathcal{Q}}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$ (another way is to do similar argumentation as in Lemma 12). Then, on the one hand, using (H-M)-(v)-(x), we have $\lim\limits_{n\rightarrow+\infty}f(s,Y^{n+1}_{s},Z^{n}_{s})=f(s,Y_{s},Z_{s})$, $d\mathbb{P}\otimes dt$-a.e. Similarly, we derive that $\lim\limits_{n\rightarrow+\infty}h(s,Y^{n+1}_{s})=h(s,Y_{s})$, $d\mathbb{P}\otimes d\kappa_{t}$-a.e. On the other hand, from (H-M)-(ix), for any $n\geq 0$ and every $\theta,\mu>0$, we infer that
$\begin{split}\mathbb{E}\left(\int_{0}^{T}f(s,Y^{n+1}_{s},Z^{n}_{s})ds\right)^{2}+\mathbb{E}\left(\int_{0}^{T}h(s,Y^{n}_{s})d\kappa_{s}\right)^{2}\leq\mathfrak{c}_{\theta,\mu,\zeta}.\end{split}$
Then, by the Lebesgue dominated convergence theorem, we can derive
$\mathbb{E}\left(\int_{0}^{T}\left(f(s,Y^{n+1}_{s},Z^{n}_{s})-f(s,Y_{s},Z_{s})\right)ds\right)^{2}+\mathbb{E}\left(\int_{0}^{T}\left(h(s,Y^{n+1}_{s})-h(s,Y_{s})\right)d\kappa_{s}\right)^{2}\xrightarrow[n\rightarrow+\infty]{}0.$ (3.44)
By the Burkholder-Davis-Gundy inequality, (H-M)-(vii), and (3.43), we have
$\begin{split}&\mathbb{E}\left[\sup_{t\in[0,T]}\left|\int_{t}^{T}\left(g(s,Y^{n}_{s},Z^{n}_{s})-g(s,Y_{s},Z_{s})\right)\overleftarrow{dB}_{s}\right|^{2}\right]\\ &\leq\mathfrak{c}\mathbb{E}\int_{0}^{T}{\Phi}^{\theta,\mu}_{s}\left|g(s,Y^{n}_{s},Z^{n}_{s})-g(s,Y_{s},Z_{s})\right|^{2}ds\\ &\leq\mathfrak{c}\left(\left\|Y^{n}-Y\right\|^{2}_{{\mathcal{H}}^{2,\mathcal{Q}}_{\theta,\mu}}+\alpha\left\{\left\|Z^{n}-Z\right\|^{2}_{{\mathcal{H}}^{2,\ell^{2}}_{\theta,\mu}}\right\}\right)\xrightarrow[n\rightarrow+\infty]{}0.\end{split}$ (3.45)
Using once more (3.43) and the Burkholder-Davis-Gundy inequality, which yields to
$\begin{split}\mathbb{E}\left[\sup_{t\in[0,T]}\left|\sum_{k=1}^{d}\int_{t}^{T}Z^{n+1,(k)}_{s}dH^{(k)}_{s}-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s}\right|^{2}\right]\leq\mathfrak{c}\left\|Z^{n+1}-Z\right\|^{2}_{{\mathcal{H}}^{2,\ell^{2}}_{\theta,\mu}}.\end{split}$ (3.46)
Next, for any $n\geq 1$, we have
[…]
From (3.43), (3.44), (3.45), (3.46), we deduce that
$K_{t}=Y_{0}-Y_{t}-\int_{0}^{t}f(s,Y_{s},Z_{s})ds-\int_{0}^{t}h(s,Y_{s})d\kappa_{s}+\int_{0}^{t}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}+\sum_{k=1}^{d}\int_{0}^{t}Z^{(k)}_{s}dH^{(k)}_{s}$.
Finally, using the same logic as in Part I of the proof of Theorem 10, we can easily see that the process $(Y,Z,K)$ solves the RGBDSDE-NL (1.5). ∎
4 Application to an obstacle problem of stochastic IPDE with nonlinear Neumann boundary conditions
This section establishes a probabilistic formula for a stochastic viscosity solution to an obstacle problem involving stochastic integral-partial differential equations with nonlinear Neumann boundary conditions (SIPDE-NBC) of parabolic type, utilizing the RGBDSDE-NL (1.5) under specific assumptions. Our technique entails initially determining the outcome for a SIPDE-NBC. The primary conclusion subsequently arises by modifying the reasons articulated in [1].
For the purpose of conducting a rigorous analysis, we assume that the inhomogeneous Lévy process $L$ lacks a Brownian component and features bounded jumps, i.e.
$L_{t}=\int_{0}^{t}{b}_{s}\,ds+\int_{0}^{t}\int_{|e|<1}e\,\tilde{\mu}(ds,de),\quad t\in[0,T].$ (4.47)
Define $\mathbf{F}^{B}\triangleq\left\{\mathcal{F}_{t,T}^{B}\right\}_{0\leq t\leq T}$. We designate by $\mathcal{M}_{0,T}^{B}$ the set of all $\mathbf{F}^{B}$-stopping times $\tau$ satisfying $0\leq\tau\leq T$, $\mathbb{P}$-almost surely. We delineate the subsequent vector spaces of functions for arbitrary Euclidean spaces $D$ and $D_{1}$:
  • $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$ is the space of all $D_{1}$-valued functions defined on $0,T]\times D$ which are $k_{1}$ times continuously differentiable in $t$ and $k_{2}$ times continuously differentiable in $x$, and $\mathcal{C}^{k_{1},k_{2}}_{b}\left([0,T]\times D;D_{1}\right)$ denotes the subspace of $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$ in which all functions have uniformly bounded partial derivatives;
  • $\mathcal{C}^{k_{1},k_{2}}(\mathcal{G},[0,T]\times D;D_{1})$ (resp. $\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathcal{G},[0,T]\times D;D_{1}\right)$), for any sub-$\sigma$-field $\mathcal{G}\subseteq\mathcal{F}_{T}^{B}$, represents the space of all $\mathcal{C}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$ (resp. $\mathcal{C}_{b}^{k_{1},k_{2}}\left([0,T]\times D;D_{1}\right)$) random variables valued in $\mathbb{R}$ that are measurable with respect to $\mathcal{G}\otimes\mathcal{B}([0,T]\times D)$.
  • $\mathcal{C}^{k_{1},k_{2}}\left(\mathbf{F}^{B},[0,T]\times D;D_{1}\right)$ (resp. $\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathbf{F}^{B},[0,T]\times D;D_{1}\right)$) is the space of all random fields $\bar{\alpha}\in\mathcal{C}^{k_{1},k_{2}}\left(\mathcal{F}_{T}^{B},[0,T]\times D;D_{1}\right)$ (resp. $\in\mathcal{C}_{b}^{k_{1},k_{2}}\left(\mathcal{F}_{T}^{B},[0,T]\times D;D_{1}\right)$), such that for fixed $x\in D$, the mapping $(t,w)\rightarrow\bar{\alpha}(t,\omega,x)$ is $\mathbf{F}^{B}$-progressively measurable.
  • $L^{p}(\mathcal{G};D)$, for any sub-$\sigma$-field $\mathcal{G}\subseteq\mathcal{F}_{T}^{B}$ and real number $p\geq 0,$ stands for all $D$-valued $\mathcal{G}$-measurable random variables $\xi$ such that $\mathbb{E}|\xi|^{p}<\infty$.
Furthermore, for $(t,x,y)\in[0,T]\times\mathbb{R}^{l}\times\mathbb{R}$, we write
$D_{x}=\left(\frac{\partial}{\partial x_{1}},\ldots,\frac{\partial}{\partial x_{l}}\right),\quad D_{xx}=\left(\partial_{x_{i}x_{j}}^{2}\right)_{i,j=1}^{l},\quad D_{y}=\frac{\partial}{\partial y},\quad D_{t}=\frac{\partial}{\partial t}.$
The meaning of $D_{xy}$ and $D_{yy}$ is then self-explanatory.
4.1 Reflected SDE driven by inhomogeneous Lèvy process
Let $G$ be an open connected bounded domain of $\mathbb{R}^{l}$ $(l\geq 1)$. Let us assume that $G$ is a smooth domain such that for a function $\Psi\in C^{2}_{b}(\mathbb{R}^{l})$, $G$ and $\partial G$ are defined as $G=\left\{x\in\mathbb{R}^{l}:\Psi(x)>0\right\}$ and $\partial G=\left\{x\in\mathbb{R}^{l}:\Psi(x)=0\right\}.$ For any $x\in\partial G$, $\nabla\Psi(x)$ represents the unit normal vector directed towards the interior of $G$. Moreover, the interior sphere condition is satisfied (see [38], pp.551), meaning there exists a $m>0$ such that for each $x\in\partial G$ and $x^{\prime}\in\bar{G}$, the following holds:
$|x^{\prime}-x|^{2}+m\langle\nabla\Psi(x),x^{\prime}-x\rangle\geq 0.$ (4.48)
For every $(t,x)\in[0,T]\times\bar{G}$, we consider the subsequent reflected stochastic differential equation driven by the inhomogeneous Lèvy process $L$ (SDEL):
[…]
where $\sigma:\mathbb{R}^{l}\to\mathbb{R}^{l}$ is globally Lipschitz, i.e. $\exists C>0,\forall x,x^{\prime}\in\mathbb{R}^{l},|\sigma(x)-\sigma(x^{\prime})|\leq C|x-x^{\prime}|,$ and satisfies the linear growth condition: $|\sigma(x)|\leq C(1+|x|)$.
First, we state some properties of the processes $(X^{t,x}_{s},\kappa^{t,x}_{s})_{s\geq 0}$ which can be found in [16] or [45].
Proposition 14.
For all $T\geq 0$, $\mu>0$ and $p\geq 2$, there exists a positive constant $C$ such that for all $t,t^{\prime}\in[0,T]$ and $x,x^{\prime}\in\bar{G}$,
  • (i)
    $\mathbb{E}\left[\underset{s\in[0,T]}{\sup}\left|X^{t,x}_{s}-X^{t^{\prime},x^{\prime}}_{s}\right|^{p}\right]+\mathbb{E}\left[\underset{s\in[0,T]}{\sup}\left|\kappa^{t,x}_{s}-\kappa^{t^{\prime},x^{\prime}}_{s}\right|^{p}\right]\leq C\left(\left|x-x^{\prime}\right|^{p}+\left|t-t^{\prime}\right|^{p/2}\right).$
  • (ii)
    $\mathbb{E}\left[\underset{s\in[0,T]}{\sup}\left|X^{t,x}_{s}\right|^{p}\right]+\mathbb{E}\left[\left|\kappa^{t,x}_{T}\right|^{p}\right]+\mathbb{E}\left[e^{\mu\kappa^{t,x}_{T}}\right]\leq C.$
Furthermore, we present the subsequent result regarding the existence and uniqueness of the solution to (4.49). The argument adheres to the same methodology as outlined in [23]. For the reader’s convenience, we delineate the principal steps of the proof.
Proposition 15.
There exists a unique solution $(X^{t,x}_{s},\kappa^{t,x}_{s})_{s\geq 0}$ for (4.49).
Proof.
The uniqueness follows directly from Proposition 14. For existence, we do the following. For the sake of brevity, we shall denote $\mathfrak{R}=\mathfrak{R}^{t,x}$ where $\mathfrak{R}=X,\kappa$. For each $n\geq 0$, define the truncated process:
$L_{t}^{n}=\int_{0}^{t}b_{s}ds+\int_{0}^{t}\int_{\frac{1}{n}<|e|<1}e\mu(ds,de)-\int_{0}^{t}\int_{\frac{1}{n}<|e|<1}eF_{s}(de)ds.$ (4.50)
Given that $L^{n}$ possesses finite variation, the Lipschitz and growth conditions are satisfied on $\sigma$, and the domain $G$ adheres to the interior sphere condition (4.48), it follows from [23] that for each $n$, a unique solution $(X^{n},\kappa^{n})$ exists for the truncated problem:
$X_{s}^{n}=x+\int_{t}^{s}\sigma(X_{r-}^{n})dL_{r}^{n}+\int_{t}^{s}\nabla\Psi(X_{r}^{n})d\kappa_{r}^{n}.$ (4.51)
Furthermore, according to Proposition 14 and Jakubowski’s compactness requirement for processes exhibiting right-continuity with left limits (see [21]), there exists a subsequence $(n_{k})_{k\geq 0}$ and a limiting process $(X,\kappa,L)$ such that:
$(X^{n_{k}},\kappa^{n_{k}},L^{n_{k}})\xrightarrow[k\to\infty]{}(X,\kappa,L).$
Finally, the limit $(X,\kappa)$ satisfies the original reflected SDEL (4.49). The convergence holds in the Skorokhod topology, and the stochastic integrals converge by continuity. ∎
4.2 Markovian generalized doubly SDE
For all $(t,x)\in[0,T]\times\bar{G}$, let $(X^{t,x}_{s},\kappa^{t,x}_{s})_{s\geq 0}$ represents the solution to the reflected SDEL (4.49). Assume that the measurable functions $f,g,h$ and $H$ satisfy:
  • $(\mathcal{H}.1)$
    $f:\Omega_{2}\times[0,T]\times\bar{G}\times\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$ satisfies for some constant $C>0$, $p\geq 2$ and some $\mathcal{F}_{t}$-measurable processes $\eta,\phi:\Omega_{2}\times[0,T]\rightarrow\mathbb{R}^{\ast}_{+}$:
    1. (i)
      $\lvert f(\omega_{2},s,x,0,0)\rvert\leq C\left(1+\left|x\right|^{p}\right);$
    2. (ii)
      $\lvert f(\omega_{2},s,X^{t,x}_{s},y,0)\rvert\leq\lvert f(\omega_{2},s,X^{t,x}_{s},0,0)\rvert+\phi_{s}|y|,$
    3. (iii)
      $(y-y^{\prime})\big(f(\omega_{2},s,X^{t,x}_{s},y,z)-f(\omega_{2},s,X^{t,x}_{s},y^{\prime},z)\big)\leq 0,$
    4. (iv)
      $\lvert f(\omega_{2},s,X^{t,x}_{s},y,z)-f(\omega_{2},s,X^{t,x}_{s},y,z^{\prime})\rvert\leq\eta_{s}\|\gamma_{s}(z-z^{\prime})\|_{\ell^{2}}$,
    5. (v)
      There exist a positive constant $C_{a}$ and $\varepsilon>0$ such that $a_{t}^{2}=\eta_{t}^{2}+\phi_{t}^{2}>\varepsilon$ and $\int_{0}^{T}a^{2}_{s}ds\leq C_{a}<\infty$.
  • $(\mathcal{H}.2)$
    $h:\Omega_{2}\times[0,T]\times\bar{G}\times\mathbb{R}\to\mathbb{R}$ satisfies for some constant $C>0$ and $p\geq 2$:
    1. (i)
      $\lvert h(s,x,0)\rvert\leq C\left(1+\left|x\right|^{p}\right),$
    2. (ii)
      $\lvert h(s,X^{t,x}_{s},y)\rvert\leq\lvert h(s,X^{t,x}_{s},0)\rvert+\zeta|y|,$
    3. (iii)
      $(y-y^{\prime})\big(h(s,X^{t,x}_{s},y)-h(s,X^{t,x}_{s},y^{\prime})\big)\leq 0.$
  • $(\mathcal{H}.3)$
    $g\in\mathcal{C}^{0,2,3}\left([0,T]\times\bar{G}\times\mathbb{R};\mathbb{R}\right)$.
  • $(\mathcal{H}.4)$
    $H:\bar{G}\to\mathbb{R}$ satisfies for some $p\geq 2$, $\lvert H(x)\rvert\leq C\left(1+\left|x\right|^{p}\right).$
As demonstrated in [12] (see Proposition 6.2), we will provide auxiliary estimations.
Proposition 16.
There exists a non-negative constant $C$ such that for all $t,t^{\prime}\in[0,T]$ and $x,x^{\prime}\in\bar{G}$,
$\mathbb{E}\left[\underset{s\in[0,T]}{\sup}e^{\mu A_{s}}\left|X^{t,x}_{s}\right|^{2}\right]\leq C\quad\text{and}\quad\mathbb{E}\left[\underset{s\in[0,T]}{\sup}e^{2\mu A_{s}}\left|X^{t,x}_{s}-X^{t^{\prime},x^{\prime}}_{s}\right|^{4}\right]\leq C.$ (4.52)
Remark 17.
Given $0,T]\times\bar{G}$ is bounded, $X^{t,x}_{s}\in\bar{G},\forall s\in[0,T]$ and the functions $f,g,h,H$ are continuous, it follows that there exists a constant $C$ such that
$\lvert f(s,X^{t,x}_{s},0,0)\rvert+\lvert h(s,X^{t,x}_{s},0)\rvert+\lvert g(s,X^{t,x}_{s},0)\rvert+\lvert H(X^{t,x}_{s})\rvert\leq C,$
and
$\mathbb{E}\left[\Phi_{T}^{\theta,\mu}\left|H(X^{t,x}_{T})\right|^{2}\right]\leq C\mathbb{E}\left[\Phi_{T}^{\theta,\mu}\left(1+\left|X^{t,x}_{T}\right|^{2}\right)\right]<\infty,$
and
$\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\frac{\left|f(s,X^{t,x}_{s},0,0)\right|^{2}}{a^{2}_{s}}ds+\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left|h(s,X^{t,x}_{s},0)\right|^{2}d\kappa^{t,x}_{s}+\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left|g(s,X^{t,x}_{s},0)\right|^{2}ds<\infty.$
Consequently, as a result of Theorem 13 below, we obtain the following outcome.
Corollary 18.
For all $(t,x)\in[0,T]\times\bar{G}$, there exists a pair of processes $(Y^{t,x}_{s},Z^{t,x}_{s})_{s\in[t,T]}$ solution to the Markovian GBDSDE-NL:
$Y^{t,x}_{s}=H(X^{t,x}_{T})+\int_{s}^{T}f(r,X^{t,x}_{r},Y^{t,x}_{r},Z^{t,x}_{r})dr+\int_{s}^{T}h(r,X^{t,x}_{r},Y^{t,x}_{r})d\kappa^{t,x}_{r}+\int_{s}^{T}g(r,X^{t,x}_{r},Y^{t,x}_{r})\overleftarrow{dB}_{r}\\ -\sum_{i=1}^{d}\int_{s}^{T}Z^{t,x(i)}_{r}dH^{(i)}_{r},~~~~t\leq s\leq T$ (4.53)
and there exists $C>0$ such that
$\mathbb{E}\left[\sup_{t\leq s\leq T}\Phi_{s}^{\theta,\mu}|Y^{t,x}_{s}|^{2}\right]+\mathbb{E}\int_{t}^{T}\Phi_{s}^{\theta,\mu}|Y^{t,x}_{s}|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{t}^{T}\Phi_{s}^{\theta,\mu}\left\|Z^{t,x}_{s}\gamma_{s}\right\|_{\ell^{2}}^{2}ds\leq C\left[1+\mathbb{E}\left(\underset{t\leq s\leq T}{\sup}\Phi_{s}^{\theta,\mu}|X^{t,x}_{s}|^{2}\right)\right]<\infty.$ (4.54)
For the remainder of this paper, we make the following assumption. Although this condition can be rigorously proven, we state it as an assumption for brevity, as the proof is lengthy and not central to our main results.
  • $(\mathcal{H}.5)$
    For any $p\geq 2$,
    $\mathbb{E}\left[\sup_{0\leq s\leq T}\left(\Phi_{s}^{\theta,\mu}\right)^{p/2}|Y^{t,x}_{s}|^{p}\right]+\mathbb{E}\left[\left(\int_{0}^{T}\Phi_{s}^{\theta,\mu}|Y^{t,x}_{s}|^{2}d\mathcal{Q}_{s}\right)^{p/2}\right]+\mathbb{E}\left[\left(\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left\|Z^{t,x}_{s}\gamma_{s}\right\|^{2}ds\right)^{p/2}\right]<\infty.$
Then, employing the same reasoning as in [37] (see also [12]), we obtain the subsequent continuity result.
Proposition 19.
Let $(Y_{s}^{t,x},Z_{s}^{t,x})$ be a solution to (4.65). The random field $(s,t,x)\rightarrow Y_{s}^{t,x},\ (s,t,x)\in[0,T]\times[0,T]\times\overline{\mathrm{G}}$ is almost surely continuous.
4.3 Stochastic viscosity solutions for SIPDE-NBC
We consider the stochastic IPDE with nonlinear Neumann boundary conditions of the form:
$\left\{\begin{array}[]{lll}\displaystyle du(t,x)+\left[\mathcal{L}_{t}u(t,x)+f\left(t,x,u(t,x),\left(u^{1}_{k}(t,x)\right)^{\infty}_{k=1}\right)\right]dt+g\left(t,x,u(t,x)\right)dB_{t}=0,&(t,x)\in[0,T[\times G\\ \\ \displaystyle u(T,x)=H(x),&x\in\bar{G}\\ \\ \displaystyle\frac{\partial u}{\partial n}(t,x)+h(t,x,u(t,x))=0,&x\in\partial G,\end{array}\right$. (4.55)
where
$\mathcal{L}_{t}u(t,x)=b_{t}.\sigma(x).D_{x}u(t,x)+\int_{\mathbb{R}}\left[u(t,x+\sigma(x)e)-u(t,x)-D_{x}u(t,x)\sigma(x)e\right]F_{t}(de)$
$u^{1}_{k}(t,x)=\int_{\mathbb{R}}\left[u(t,x+\sigma(x)e)-u(t,x)-D_{x}u(t,x)e\right]p_{k}(e)F_{t}(de)$
and, for every $x\in\partial G$,
${\frac{\partial\varphi}{\partial n}=\sum_{i=1}^{l}\frac{\partial\Psi}{\partial x_{i}}(x)\frac{\partial\varphi}{\partial x_{i}}(t,x).}$
We now introduce the concept of stochastic viscosity solution for the SIPDE $(f,g,h)$. We are inspired by the work of Buckdahn and Ma [6], and we refer to their paper for a clear exposition on this topic. We use part of their nomenclature and follow the lines of their arguments to arrive at our major conclusion.
We shall utilize the stochastic flow $\hat{\chi}(t,x,y)\in\mathcal{C}(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{l}\times\mathbb{R};\mathbb{R})$, which is the only solution to the SDE that looks like this in Stratonovich form:
$\hat{\chi}(t,x,y)=y+\int_{0}^{t}g(s,x,\hat{\chi}(s,x,y))\circ dB_{s},\quad t\geq 0.$
Under $(\mathcal{H}.3)$, the mapping $y\mapsto\hat{\chi}(t,x,y)$ constitutes a diffeomorphism for every $(t,x)$, $\mathbb{P}$-almost surely (see [42]). Let the $y$-inverse of $\hat{\chi}(t,x,y)$ be represented as $\tilde{\pi}(t,x,y)$. Consequently, given that $\tilde{\pi}(t,x,\hat{\chi}(t,x,y))=y$, it can be demonstrated (see [6] for instance)
$\tilde{\pi}(t,x,y)=y-\int_{0}^{t}D_{y}\tilde{\pi}(s,x,y)g(s,x,y)\circ dB_{s},$
where the stochastic integrals must be read in the Stratonovich sense.
We will now introduce the process $\chi\in\mathcal{C}(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{n}\times\mathbb{R};\mathbb{R})$ as the solution to the equation
$\chi(t,x,y)=y+\int_{t}^{T}g(s,x,\chi(s,x,y))\circ dB_{s},\quad 0\leq t\leq T.$ (4.56)
It is observed that, owing to the direction of the Itô integral, equation (4.56) must be seen as progressing from $T$ to $t$ (i.e., $y$ should be regarded as the beginning value). Consequently, the mapping $y\mapsto\chi(s,x,y)$ will exhibit same regularity characteristics as the mapping $y\mapsto\hat{\chi}(s,x,y)$ for all $(s,x)\in[t,T]\times\mathbb{R}^{n}$, $\mathbb{P}$-almost surely. Thus, if we indicate its $y$-inverse by $\pi$, we obtain
$\pi(t,x,y)=y-\int_{t}^{T}D_{y}\pi(s,x,y)g(s,x,y)\circ dB_{s}.$
To simplify the notation we write:
$\mathcal{L}^{t}_{f,g}(\varphi(t,x))=-\mathcal{L}_{t}\varphi(t,x)-f\left(t,x,\varphi(t,x),\left(\varphi^{1}_{k}(t,x)\right)^{\infty}_{k=1}\right)+\frac{1}{2}g.D_{y}g(t,x,\varphi(t,x)).$
We now introduce the notion of a stochastic viscosity solution of (4.55) as follows.
Definition 20.
  • (1).
    A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\overline{\mathrm{G}};\mathbb{R}\right)$ is called a stochastic viscosity subsolution (resp. supersolution) of (4.55) if $u(T,x)\leq H(x)$ (resp. $\geq$), for all $x\in\overline{\mathrm{G}}$, and if for any stopping time $\tau\in\mathcal{M}_{0,T}^{B}$, any state variable $\xi\in L^{0}\left(\mathcal{F}_{\tau}^{B},G\right)$, and any random field $\varphi\in\mathcal{C}^{1,2}\left(\mathcal{F}_{\tau}^{B},[0,T]\times G;\mathbb{R}\right)$, with the property that for $\mathbb{P}^{2}$-almost all $\omega_{2}\in\{0<\tau<T\}$ the inequality
    $u(t,\omega_{2},x)-\chi(t,\omega_{2},x,\varphi(t,x))\leq 0=u(\tau(\omega_{2}),\xi(\omega_{2}))-\chi(\tau(\omega_{2}),\xi(\omega_{2}),\varphi(\tau(\omega_{2}),\xi(\omega_{2})))\ (\text{resp.}\ \geq)$
    is fulfilled for all $(t,x)$ in some neighbourhood of $(\tau(\omega_{2}),\xi(\omega_{2})$), the following conditions are satisfied:
    1. (a)
      On the event $\{0<\tau<T\}$ the inequality
      $\mathcal{L}^{t}_{f,g}(\psi(\tau,\xi))-D_{y}\chi(\tau,\xi,\psi(\tau,\xi))D_{t}\varphi(\tau,\xi)\leq 0.\ \ (\text{resp.}\ \geq)$ (4.57)
      holds $\mathbb{P}^{2}$-almost surely, where $\psi(t,x)\triangleq\chi(t,x,\varphi(t,x))$.
    2. (b)
      On the event $\{0<\tau<T\}\cap\{\xi\in\partial\mathrm{G}\}$ the inequality
      $\min\left[\mathcal{L}^{t}_{f,g}(\psi(\tau,\xi))-D_{y}\chi(\tau,\xi,\varphi(\tau,\xi))D_{t}\varphi(\tau,\xi)-\frac{\partial\psi}{\partial n}(\tau,\xi)-h(\tau,\xi,\psi(\tau,\xi))\right]\leq 0.\ \ (\text{resp.}\ \geq)$ (4.58)
      holds $\mathbb{P}$-almost surely with $\psi(t,x)\triangleq\chi(t,x,\varphi(t,x))$.
  • (2)
    A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\overline{\mathrm{G}};\mathbb{R}\right)$ is called a stochastic viscosity solution of (4.55) if it is both a stochastic viscosity subsolution and a supersolution.
Remark 21.
Note that if $f,h$ are deterministic and $g\equiv 0$, then Definition 20 coincides with the deterministic case (see, e.g., [12]).
Now, let us recall the concept of a random viscosity solution, which will serve as a connection between the stochastic viscosity solution and its deterministic equivalent.
Definition 22.
A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{l};\mathbb{R}\right)$ is called an $\omega$-wise viscosity solution if for $\mathbb{P}$-almost all $\omega\in\Omega,u(\omega,\cdot,\cdot)$ is a (deterministic) viscosity solution of the $\operatorname{SIPDE}(f,0,h)$.
To proceed, let us present the following lemma adapted from [27].
Lemma 23.
Let $\bar{h}:\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ be a $\mathcal{P}\otimes\mathcal{B}_{\mathbb{R}}$-measurable function such that
$|\bar{h}(s,y)|\leq\tilde{a}_{s}(e^{2}\wedge|e|)\quad\text{a.s.},$
where $(\tilde{a}_{s})_{s\geq 0}$ is a positive predictable process with $\mathbb{E}\left[\int_{0}^{T}\tilde{a}_{s}^{2}ds\right]<\infty$. Then
$\sum_{t\leq s\leq T}\bar{h}(s,\Delta L_{s})=\sum_{i=1}^{d}\int_{t}^{T}\langle\bar{h}(s,\cdot),p_{i}\rangle_{L^{2}(\nu)}dH_{s}^{(i)}+\int_{t}^{T}\int_{\mathbb{R}}\bar{h}(s,e)F_{s}(de)ds,$
where
$\langle\bar{h}(s,\cdot),p_{i}\rangle_{L^{2}(\nu)}=\int_{\mathbb{R}}\bar{h}(s,e)p_{i}(e)F_{s}(de)ds.$
[…]
Remark 24.
Recall that under assumption $(\mathcal{H}.3)$, the random field $\chi$ belongs to $C^{0,2,2}\left(\mathbf{F}^{B}\right$., $0,T]\times\mathbb{R}^{n}\times\mathbb{R};\mathbb{R}$) and hence, it is true for $\pi$. Subsequently, by using the transformation $\psi(t,x)=\chi(t,x,\varphi(t,x))$, we derive
$D_{x}\psi$ $=D_{x}\chi+D_{y}\chi D_{x}\varphi,$
$D_{xx}\psi$ $=D_{xx}\chi+2\left(D_{xy}\chi\right)\left(D_{x}\varphi\right)^{*}+\left(D_{yy}\chi\right)\left(D_{x}\varphi\right)\left(D_{x}\varphi\right)^{*}+\left(D_{y}\chi\right)\left(D_{xx}\varphi\right).$
Moreover, since for all $(t,x,y)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}$ the equality $\pi(t,x,\chi(t,x,y))=y$ holds $\mathbb{P}$ almost surely, we also have
$D_{x}\pi+D_{y}\pi D_{x}\chi$ $=0,$
$D_{y}\pi D_{y}\chi$ $=1,$
$D_{xx}\pi+2\left(D_{xy}\pi\right)\left(D_{x}\chi\right)^{*}+\left(D_{yy}\pi\right)\left(D_{x}\chi\right)\left(D_{x}\chi\right)^{*}+\left(D_{y}\pi\right)\left(D_{xx}\chi\right)$ $=0$
$\left(D_{xy}\pi\right)\left(D_{y}\chi\right)+\left(D_{yy}\pi\right)\left(D_{x}\chi\right)\left(D_{y}\chi\right)+\left(D_{y}\pi\right)\left(D_{xy}\chi\right)$ $=0,$
$\left(D_{yy}\pi\right)\left(D_{y}\chi\right)^{2}+\left(D_{y}\pi\right)\left(D_{yy}\chi\right)$ $=0,$
where all the derivatives of the random field $\pi(\cdot,\cdot,\cdot)$ are evaluated at $(t,x,\chi(t,x,y)$), whereas all derivatives of $\chi(\cdot,\cdot,\cdot)$ are evaluated at $(t,x,y)$.
Proposition 25.
[…]
Proof.
[…]
$v(t,x)-\varphi(t,x)\leq 0=v(\tau(\omega),\xi(\omega))-\varphi(\tau(\omega),\xi(\omega))$
holds for all $(t,x)$ in some neighbourhood $\mathcal{V}(\omega,\tau(\omega),\xi(\omega))$ of $(\tau(\omega),\xi(\omega))$. Next we put $\psi(t,x)=\chi(t,x,\varphi(t,x))$. Since the mapping $y\mapsto\chi(t,x,y)$ is strictly increasing, for all $(t,x)\in$ $\mathcal{V}(\tau,\xi)$ we have that
$u(t,x)-\psi(t,x)$ $=\chi(t,x,v(t,x))-\chi(t,x,\varphi(t,x))$
$\leq 0=\chi(\tau,\xi,v(\tau,\xi))-\chi(\tau,\xi,\varphi(\tau,\xi))$
$=u(\tau,\xi)-\psi(\tau,\xi)$
holds $\mathbb{P}$-almost surely on $\{0<\tau<T\}$. Moreover, since $u$ is a stochastic viscosity subsolution of the $\operatorname{SIPDE}(f,g,h)$, the inequality
$\mathcal{L}^{t}_{f,g}(\psi(\tau,\xi))-D_{y}\chi(\tau,\xi,\varphi(\tau,\xi))D_{t}\varphi(\tau,\xi)\leq 0,$ (4.59)
holds $\mathbb{P}$-almost everything on the event $\{0<\tau<T\}$. On the other hand, from Remark 24 it follows that
$\mathcal{L}_{t}\psi(t,x)$ $=b_{t}.\sigma(x).D_{x}\psi(t,x)+\int_{\mathbb{R}}\left[\psi(t,x+\sigma(x)e)-\psi(t,x)-D_{x}\psi(t,x)\sigma(x)e\right]F_{t}(de)$
$=\mathcal{L}_{t}^{x}\chi(t,x,\varphi(t,x))+D_{y}\chi\cdot\mathcal{L}_{t}\varphi(t,x)$
$+\int_{\mathbb{R}}\Big[\chi(t,x+\sigma(x)e,\varphi(t,x+\sigma(x)e))-\chi(t,x+\sigma(x)e,\varphi(t,x))-D_{y}\chi\cdot D_{x}\varphi\cdot\sigma(x)e\Big]F_{t}(de).$
Then, if we define the random field $\widetilde{f}$ by
$\tilde{f}(t,x,y,z)$ $=\frac{1}{D_{y}\chi}\Bigg[f\bigg(t,x,\chi,\left(\tilde{\psi}^{1}_{k}\right)_{k=1}^{\infty}\bigg)-\frac{1}{2}gD_{y}g+\mathcal{L}_{t}^{x}\chi$
$+\int_{\mathbb{R}}\Bigl[\chi\bigl(t,x+\sigma(x)e,\,\varphi(t,x+\sigma(x)e)\bigr)-\chi-\bigl(D_{x}\chi+D_{y}\chi\cdot z\bigr)\sigma(x)e\Bigr]F_{t}(de)\Bigg]$ (4.60)
where $\chi=\chi(t,x,y)$ and $\tilde{\psi}^{1}_{k}$ is defined as follows
$\tilde{\psi}^{1}_{k}(t,x)=\int_{\mathbb{R}}\left[\chi(t,x+\sigma e,\varphi(t,x+\sigma e))-\chi-(D_{x}\chi+D_{y}\chi D_{x}\varphi)\cdot e\right]p_{k}(e)F_{t}(de)$
we obtain
$D_{y}\pi(t,x,\psi(t,x))\mathcal{L}^{t}_{f,g}(\psi(t,x))=\mathcal{L}^{t}_{\tilde{f},0}(\varphi(t,x)).$
Consequently, (4.59) becomes
$\mathcal{L}^{t}_{\tilde{f},0}(\varphi(\tau,\xi))-D_{t}\varphi(\tau,\xi)\leq 0,$ (4.61)
and hence the Doss-Sussman transformation converts an SIPDE of the form $(f,g,h)$ to one of the form $(f,0,h)$, provided a similar transformation for the random field $h$ also works for the inequality in (4.57). This establishes part (a) of Definition 20.
To establish part (b) in Definition 20, we notice that for every $(t,x)\in[0,T]\times\partial\mathrm{G}$, the subsequent series of equalities is valid:
$\frac{\partial\psi}{\partial n}(t,x)$ $=\left\langle D_{x}\psi(t,x),\nabla\Psi(x)\right\rangle$
$=\left\langle D_{x}\chi(t,x,\varphi(t,x)),\nabla\Psi(x)\right\rangle+D_{y}\chi(t,x,\varphi(t,x))\left\langle D_{x}\varphi(t,x),\nabla\psi(x)\right\rangle$
$=\left\langle D_{x}\chi(t,x,\varphi(t,x)),\nabla\Psi(x)\right\rangle+D_{y}\chi(t,x,\varphi(t,x))\frac{\partial\varphi}{\partial n}(t,x).$
Hence,
$\frac{\partial\psi}{\partial n}(\tau,\xi)+h(\tau,\xi,\psi(\tau,\xi))=$ $D_{y}\chi(\tau,\xi,\varphi(\tau,\xi))\frac{\partial\varphi}{\partial n}(\tau,\xi)+\left\langle D_{x}\chi(\tau,\xi,\varphi(\tau,\xi)),\nabla\Psi(x)\right\rangle+h(\tau,\xi,\eta(\tau,\xi,\varphi(\tau,\xi)))$
$=$ $D_{y}\chi(\tau,\xi,\varphi(\tau,\xi))\left(\frac{\partial\varphi}{\partial n}(\tau,\xi)+\widetilde{h}(\tau,\xi,\varphi(\tau,\xi))\right)$
where
$\widetilde{h}(t,x,y)=\frac{1}{D_{y}\chi(t,x,y)}\left(h(t,x,\chi(t,x,y))+\left\langle D_{x}\chi(t,x,y),\nabla\Psi(x)\right\rangle\right).$ (4.62)
Since $D_{y}\chi(t,x,y)>0$, we obtain, $\mathbb{P}$-almost surely on the event $\{0<\tau<T\}\cap\{\xi\in\partial\mathrm{G}\}$, the inequality
$\min\left[\mathcal{L}^{t}_{\tilde{f},0}(\varphi(\tau,\xi))-D_{t}\varphi(\tau,\xi),-\frac{\partial\varphi}{\partial n}(\tau,\xi)-\widetilde{h}(\tau,\xi,\psi(\tau,\xi))\right]\leq 0.$ (4.63)
By integrating (4.61) and (4.63), we establish that the random field $v$ is a stochastic viscosity subsolution of $\operatorname{SIPDE}(\widetilde{f},0,\widetilde{h}), the$refore concluding the proof of Proposition 25. ∎
To this end, we need the following result which is proved in Buckdahn and Ma [2]:
Proposition 26.
Assume $\left(\mathcal{H}.{5}\right)$ holds. Let $\chi$ be the unique solution to $\operatorname{SDE}$ (4.56) and $\pi$ be the $y$-inverse of $\chi$. Then, there exists a constant $C>0$, depending only on the bound of $g$ and its partial derivatives, such that for $\zeta=\chi,\pi$, the following inequalities hold $\mathbb{P}$-almost surely for all $(t,x,y)\in[0,T]\times\mathbb{R}^{l}\times\mathbb{R}$:
$|\zeta(t,x,y)|\leq|y|+C\left|B_{t}\right|$
$\left|D_{x}\zeta\right|,\left|D_{y}\zeta\right|,\left|D_{xx}\zeta\right|,\left|D_{xy}\zeta\right|,\left|D_{yy}\zeta\right|\leq C\exp\left\{C\left|B_{t}\right|\right\}$
Here all the derivatives are evaluated at $(t,x,y)$.
Next, for $t\in[0,T]$ and $x\in\bar{G}$, let us define the following processes:
$U_{s}^{t,x}=\chi\left(s,X_{s}^{t,x},Y_{s}^{t,x}\right),\quad 0\leq t\leq s\leq T$
${V_{s}^{t,x}=\int_{\mathbb{R}}\Bigl[\chi\bigl(t,X_{s}^{t,x}+\sigma(X_{s}^{t,x})e,\,Y_{s}^{t,x}\bigr)-\chi-\bigl(D_{x}\chi+D_{y}\chi\cdot Z^{t,x}_{s}\bigr)\sigma(X_{s}^{t,x})e\Bigr]F_{s}(de)}$
Then, from Proposition 26, we obtain
$\left(\left(U_{s}^{t,x},V_{s}^{t,x}\right),(s,x)\in[0,T]\times\overline{\mathrm{G}}\right)\in\mathcal{S}^{2}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$
Theorem 27.
For each $(t,x)\in[0,T]\times\overline{\mathrm{G}}$, the process $\left(\left(U_{s}^{t,x},V_{s}^{t,x}\right),(s,x)\in[0,T]\times\overline{\mathrm{G}}\right)$ is the unique solution to the following generalized BSDE:
$U_{s}^{t,x}=H\left(X_{T}^{t,x}\right)+\int_{s}^{T}\tilde{f}\left(t,X_{r}^{t,x},U_{r}^{t,x},V_{r}^{t,x}\right)\mathrm{d}r+\int_{s}^{T}\tilde{h}\left(r,X_{r}^{t,x},U_{r}^{t,x}\right)\mathrm{d}k_{r}^{t,x}-\sum_{i=1}^{d}\int_{s}^{T}V^{t,x(i)}_{r}dH^{(i)}_{r}$, (4.64)
where $\tilde{f}$ and $\tilde{h}$ are given by (4.3) and (4.62).
Here, the argument adheres to the same methodology as outlined in [4] (Theorem 3.4), together with Lemma 6. Therefore, we omit the proof.
Now, define for each $(t,x)\in[0,T]\times\overline{\mathrm{G}}$ the random fields $u$ and $v$ by $u(t,x)=Y_{t}$ and $v(t,x)=U_{t}$, where $(Y,Z)$ and $(U,V)$ are the solutions to the equations (4.53) and (4.64), respectively. Then, we have
$u(\omega,t,x)=\chi(\omega,t,x,v(t,\omega,x)),\quad v(\omega,t,x)=\pi(\omega,t,x,u(t,\omega,x)).$
Subsequently, we obtain the following result, whose proof adheres to the same reasoning as shown in [4].
Theorem 28.
The random field $u(t,x)=Y^{t,x}_{t}$ is a stochastic viscosity solution to the $\operatorname{SIPDE}(f,g,h)$.
4.4 The obstacle problem of SIPDE-NBC
In this final segment, we report our primary finding, extending the previous results to formulate a probabilistic formula for an obstacle problem of SIPDE-NBC.
Initially, we shall incorporate the subsequent assumption on the barrier.
  • $(\mathcal{H}.6)$
    $\mathcal{S}:[0,T]\times\bar{G}\to\mathbb{R}$ satisfies for all $s\in[0,T]$ and $p\geq 2$:
    1. (i)
      $\lvert\mathcal{S}(s,x)\rvert\leq C\left(1+\left|x\right|^{p}\right),$
    2. (ii)
      $\mathcal{S}\in C^{1,2}~~\text{such that}\ \mathcal{S}(T,x)\leq H(x),~~\forall x\in\bar{G}.$
As in Remark 24, we have
$\lvert\mathcal{S}(s,X^{t,x}_{s})\rvert\leq C\ \ \text{and}\ \ \mathbb{E}\left[\sup_{0\leq s\leq T}\Phi_{s}^{\theta,\mu}\left|\mathcal{S}(s,X^{t,x}_{s})\right|^{2}\right]\leq C\mathbb{E}\left[\Phi_{s}^{\theta,\mu}\left(1+\left|X^{t,x}_{s}\right|^{2}\right)\right]<\infty.$
Therefore, by Theorem 13 below we have the following result.
Corollary 29.
For all $(t,x)\in[0,T]\times\bar{G}$, there exists a triplet of processes $(Y^{t,x}_{s},Z^{t,x}_{s},K^{t,x}_{s})_{s\in[t,T]}$ solution to the Markovian Reflected GBDSDE-NL:
$Y^{t,x}_{s}=H(X^{t,x}_{T})+\int_{s}^{T}f(r,X^{t,x}_{r},Y^{t,x}_{r},Z^{t,x}_{r})dr+\int_{s}^{T}h(r,X^{t,x}_{r},Y^{t,x}_{r})d\kappa^{t,x}_{r}+\int_{s}^{T}g(r,X^{t,x}_{r},Y^{t,x}_{r})\overleftarrow{dB}_{r}\\ +K^{t,x}_{T}-K^{t,x}_{s}-\sum_{i=1}^{d}\int_{s}^{T}Z^{t,x(i)}_{r}dH^{(i)}_{r},~~~~t\leq s\leq T$ (4.65)
and there exists $C>0$ such that
$\mathbb{E}\left[\sup_{t\leq s\leq T}\Phi_{s}^{\theta,\mu}|Y^{t,x}_{s}|^{2}\right]+\mathbb{E}\int_{t}^{T}\Phi_{s}^{\theta,\mu}|Y^{t,x}_{s}|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{t}^{T}\Phi_{s}^{\theta,\mu}\left\|Z^{t,x}_{s}\gamma_{s}\right\|_{\ell^{2}}^{2}ds+\mathbb{E}\left|K^{t,x}_{T}-K^{t,x}_{t}\right|^{2}\\ \leq C\left[1+\mathbb{E}\left(\underset{t\leq s\leq T}{\sup}\Phi_{s}^{\theta,\mu}|X^{t,x}_{s}|^{2}\right)\right]<\infty.$ (4.66)
Using the same arguments as in [37], we get the following continuity result:
Proposition 30.
Let $(Y_{s}^{t,x},Z_{s}^{t,x})$ be a solution to (4.65). Then, the random field $(s,t,x)\rightarrow Y_{s}^{t,x},\ (s,t,x)\in[0,T]\times[0,T]\times\overline{\mathrm{G}}$ is almost surely continuous.
Utilizing the results from Section 4.3 and employing the methodology outlined in [1] (Theorem 4.2, pp.1168), we derive:
Theorem 31.
Under conditions $(\mathcal{H}.1)$$(\mathcal{H}.6)$, the random field $u(\omega,t,x)=Y^{t,x}_{t}$ is stochastic viscosity solution for the subsequent obstacle problem of SIPDE-NBC
$\left\{\begin{array}[]{lll}\displaystyle{\big(u(t,x)-\mathcal{S}(t,x)\big)\wedge\left(-\frac{\partial u}{\partial t}(t,x)-\mathcal{L}_{t}u(t,x)-f\left(t,x,u(t,x),\left(u^{1}_{k}(t,x)\right)^{\infty}_{k=1}\right)\right)}\\ \hskip 256.0748pt\displaystyle+g\left(t,x,u(t,x)\right)dB_{t}=0,&(t,x)\in[0,T[\times G\\ \\ \displaystyle u(T,x)=H(x),&x\in\bar{G}\\ \displaystyle\frac{\partial u}{\partial n}(t,x)+h(t,x,u(t,x))=0,&x\in\partial G,\end{array}\right$.
Appendix A Special generalized backward doubly SDEs with Lipschitz drivers
The existence and uniqueness results for generalized BDSDEs with a non-homogeneous Lévy process (GBDSDE-NL) are provided in this section for the specific case where the coefficients $f$ and $g$ are Lipschitz and depend only on $y$, while the driver $g$ does not depend on the variables $(y,z)$. We first consider the case of standard generalized BDSDE-NL without reflection. Then, using these results, we extend the analysis to the reflected case with a continuous obstacle $S$.
We denote $\mathfrak{f}(t,y):=f(t,y,z)$ and $\mathfrak{g}(t):=g(t,y,z)$ for any $(t,y,z)\in[0,T]\times\mathbb{R}\times\ell^{2}$. Toward the end of this section, the notations from the previously mentioned assumption (H-M) are used, along with the same satisfied properties.
A.1 Classical generalized backward doubly SDEs
Consider the following GBDSDE-NL:
$Y_{t}=\xi+\int_{t}^{T}\mathfrak{f}(s,Y_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}+\int_{t}^{T}\mathfrak{g}(s)\overleftarrow{dB}_{s}-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s},\quad t\in[0,T].$ (A.67)
Theorem 32 (Existence and uniqueness).
Assume that
(H-LP)
  • There exists a constant $\mathfrak{b}>0$ such that, for any $t\in[0,T]$ and $y,y^{\prime}\in\mathbb{R}$, we have
    $\left|\mathfrak{f}(t,y)-\mathfrak{f}(t,y^{\prime})\right|+\left|h(t,y)-h(t,y^{\prime})\right|\leq\mathfrak{b}\left|y-y^{\prime}\right|.$
  • For any $t\in[0,T]$, we have $\left|\mathfrak{f}(t,0)\right|\leq\varphi_{t}$ and $\left|h(t,0)\right|\leq\psi_{t}$.
  • For all $\mu>0$:
    • $\mathbb{E}\left(e^{\mu\kappa_{T}}|\xi|^{2}\right)<+\infty.$
    • $\mathbb{E}\left(\int_{0}^{T}e^{\mu\kappa_{t}}\left(\left|\frac{\varphi_{t}}{a_{t}}\right|^{2}+\left|\mathfrak{g}(t)\right|^{2}\right)dt\right)+\mathbb{E}\left(\int_{0}^{T}e^{\mu\kappa_{t}}\left|\psi_{t}\right|^{2}d\kappa_{t}\right)<+\infty$.
Then the BDSDE-NL (A.67) admits a unique solution $(Y,Z)\in\left(\mathcal{S}^{2}_{\mu}\cap\mathcal{H}^{2,A}_{\mu}\right)\times\mathcal{H}^{2,\ell^{2}}_{\mu}$.
Proof.
By adopting a similar approach to that used in [20, Theorem 3.2], along with the predictable representation theorem for non-homogeneous Lévy processes provided in [8, Theorem 2.4] (see also [8, Section 2.3. Example 3]), we can establish the existence of a pair $(Y,Z)\in\mathcal{S}^{2}\times\mathcal{H}^{2,\ell^{2}}$ that satisfies the BDSDE-NL (A.67). The integrability condition is verified by following the same computations as in Proposition 8, where it suffices to prove the result in the case where $\kappa_{T}$ is a bounded random variable, and subsequently apply Fatou’s Lemma (see, for instance, [15, Proposition 1] for a similar argument). Finally, the uniqueness result can be deduced in a similar way, following the argument in Proposition 9. ∎
Theorem 33 (Comparison theorem).
Assume that condition (H-LP) holds. Let $(Y^{i},Z^{i})$ be a solution to the BDSDE-NL (A.67) associated with the parameters $(\xi^{i},\mathfrak{f}^{i},h,\kappa,\mathfrak{g})$ for $i=1,2$. If $\xi^{1}\leq\xi^{2}$ and $\mathfrak{f}^{1}(t,y)\leq\mathfrak{f}^{2}(t,y)$ for all $(t,y)\in[0,T]\times\mathbb{R}$, then $Y^{1}_{t}\leq Y^{2}_{t}$ a.s. for all $t\in[0,T]$.
Proof.
The proof follows the same reasoning as in the proof of [1, Lemma 3.1]. Hence, we omit the details. ∎
A.2 Special generalized reflected backward doubly SDEs
We consider the following version of the RGBDSDE-NL (1.5):
$\left\{\begin{split}\text{(i)}&~Y_{t}=\xi+\int_{t}^{T}\mathfrak{f}(s,Y_{s})ds+\int_{t}^{T}h(s,Y_{s})d\kappa_{s}+\int_{t}^{T}\mathfrak{g}(s)\overleftarrow{dB}_{s}+\left(K_{T}-K_{t}\right)-\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s},\\ \text{(ii)}&~Y_{t}\geq S_{t},~t\in[0,T]~\mbox{ and }~\int_{0}^{T}\left(Y_{s}-S_{s}\right)dK_{s}=0.\end{split}\right$. (A.68)
Theorem 34.
Assume that the data $\left(\xi,\mathfrak{f},h,\kappa,\mathfrak{g}\right)$ satisfy assumption (H-LP) and that the obstacle $S$ satisfies $\mathbb{E}\sup_{t\in[0,T]}\left|e^{\mu}S^{+}_{t}\right|^{2}<+\infty$ for any $\mu>0$. Then the RGBDSDE-NL (A.68) admits a unique solution $\left(Y,Z,K\right)\in\left(\mathcal{S}^{2}_{\mu}\cap\mathcal{H}^{2,A}_{\mu}\right)\times\mathcal{H}^{2,\ell^{2}}_{\mu}\times\mathcal{A}^{2}$.
Proof.
The proof is quite standard, using Theorems 32 and $3$, and following a similar argument as in [1, Proposition 3.2]. The integrability property follows from Proposition 8, in a manner similar to that established in Theorem 32. ∎
Disclosure statement
No potential conflict of interest was reported by the authors.